### v1

# Active Hypothesis Testing under Computational Budgets with Applications to GWAS and LLM

Abstract
In large-scale hypothesis testing, computing exact $p$-values or $e$-values is often resource-intensive, creating a need for budget-aware inferential methods. We propose a general framework for active hypothesis testing that leverages inexpensive auxiliary statistics to allocate a global computational budget. For each hypothesis, our data-adaptive procedure probabilistically decides whether to compute the exact test statistic or a transformed proxy, guaranteeing a valid $p$-value or $e$-value while satisfying the budget constraint in expectation. Theoretical guarantees are established for our constructions, showing that the procedure achieves optimality for $e$-values and for $p$-values under independence, and admissibility for $p$-values under general dependence. Empirical results from simulations and two real-world applications, including a large-scale genome-wide association study (GWAS) and a clinical prediction task leveraging large language models (LLM), demonstrate that our framework improves statistical efficiency under fixed resource limits.
Keywords: Active Learning, Budget-aware Inference, Computational Constraints, $e$-value, Multiple Testing, $p$-value
1 Introduction
The $p$-value and $e$-value (Vovk and Wang,, 2021; Ren and Barber,, 2023; Ramdas and Wang,, 2024) are fundamental tools in statistical inference for quantifying evidence against a null hypothesis. While essential, their exact computation can be prohibitively expensive due to costly experimental procedures or substantial computational demands. This challenge creates a need for inferential methods that operate within a fixed budget. We propose a general framework for active hypothesis testing that addresses this problem directly. Our approach leverages inexpensive and readily available auxiliary statistics, which are derived from cheaper data sources, prior knowledge, or predictive models, to manage a global computational budget. For each hypothesis, a data-adaptive procedure probabilistically decides whether to compute the resource-intensive “gold-standard” statistic. When the exact statistic is not computed, a transformed version of the auxiliary statistic is used in its place, ensuring a valid test statistic for every hypothesis. This framework finds wide applicability across various domains, as illustrated by the following examples.
  1. (a)
    Powerful Prediction Model. Consider a setting where a large, pre-trained prediction model is available to forecast an outcome of interest (Angelopoulos et al.,, 2023; Motwani and Witten,, 2023; Zrnic and Candès,, 2024; Kluger et al.,, 2025; Ji et al.,, 2025). The exact $e$-value or $p$-value requires observing the actual outcomes, which may be costly or delayed. A proxy statistic can be rapidly computed by using the model’s predictions in place of the true outcomes. In this setting, the proxy statistic acts as an auxiliary statistic to guide whether observing the true outcomes is worth the cost or delay.
  2. (b)
    Costly vs. Noisy Measurements. In many scientific and industrial domains, a precise measurement is destructive, time-consuming, or financially expensive, while a cheaper but noisier measurement is often available from alternative sensors (Carroll et al.,, 1995; Fuller,, 2009; Grace et al.,, 2021; Dunbar et al.,, 2022). For instance, a full genetic assay is costly, but a simple biomarker measurement is not. A valid $p$-value or $e$-value can only be derived from the precise measurement, while the noisy data provides an informative, yet potentially biased, proxy. Our framework formally navigates this trade-off, using the noisy measurement as an informative guide to determine when the budget should be spent on the definitive, precise measurement.
  3. (c)
    Multi-view Learning and Complementary Signals. In many applications, data provide multiple distinct “views” of the same underlying phenomenon, where different perspectives can offer complementary information (Zhang et al.,, 2011; Sun,, 2013; Zhao et al.,, 2017). A common setting involves two complementary data views. The first view, e.g., genomic or genotype data, is expensive to collect but supports valid computation of $p$-values or $e$-values. The second view, such as routine clinical measurements or gene expression data, is inexpensive to obtain but its null distribution is unknown, making it unsuitable for direct inferential procedures. Despite this limitation, the second view can provide substantial predictive signal. Our framework synthesizes these complementary sources of information by using the cheap data view as a powerful proxy to strategically allocate the budget for computations on the resource-intensive view.
A central challenge addressed in this work is the efficient integration of two types of information: a costly but statistically valid test statistic, and an auxiliary statistic that is inexpensive to obtain but may be less reliable. Formally, we consider a setting with $N$ hypotheses, each associated with a resource-intensive test statistic that yields valid inference and a cheap auxiliary statistic that may be unreliable. Our objective is to construct a valid test statistic for every hypothesis while ensuring that the expected number of costly computations does not exceed a predetermined global budget.
1.1 Related Work
Prior work on incorporating auxiliary statistics in hypothesis testing has largely focused on improving statistical power, with limited consideration of computational budget constraints. These approaches typically leverage side information to prioritize hypotheses, and are often implemented through weighted multiple testing procedures, which can be interpreted either as re-weighting the $p$-values (Genovese et al.,, 2006; Ignatiadis et al.,, 2016; Liu et al.,, 2016; Barber and Ramdas,, 2017; Xia et al.,, 2020; Cai et al.,, 2022) or, equivalently, as adaptively adjusting the rejection thresholds (Lei and Fithian,, 2018; Zhang et al.,, 2019; Li and Barber,, 2019; Chao and Fithian,, 2021; Freestone et al.,, 2024). Although these approaches improve power, they rely on the assumption that an exact $p$-value is available for every hypothesis and therefore do not tackle the fundamental challenge of high computational or experimental cost. Nevertheless, the weights or prioritization scores generated by some of these approaches can be leveraged as auxiliary statistics within our framework to inform efficient budget allocation. Similarly, two-stage multiple testing procedures use an inexpensive screening stage to filter out unpromising hypotheses (Zehetmayer et al.,, 2005; Aoshima and Yata,, 2011). These methods, however, typically rely on a hard selection rule: hypotheses that fail the initial screening are discarded, and no formal inferential statements are made for them.
A second line of research, which inspires our approach, is active learning, where information is queried selectively to improve efficiency (Cohn et al.,, 1996; Settles,, 2009; Sener and Savarese,, 2018; Ren et al.,, 2021). However, our goal is fundamentally different. The active learning literature, including recent work on acquiring gold-standard labels for statistical inference (Zhang et al.,, 2021; Zrnic and Candès,, 2024; Cook et al.,, 2024), has primarily focused on optimizing the collection of labeled data for parameter estimation or model training. While conceptually related, these methods aim to improve the efficiency of data collection, whereas our work focuses on the dynamic decision of whether to compute a test statistic itself.
The work most closely related to ours is the recently proposed proxy computing framework of Xu et al., 2025b , which introduces an innovative approach to reducing the computational burden of hypothesis testing by probabilistically deciding whether to compute an exact statistic. Our work builds upon this foundation by introducing a key extension: we develop a flexible, global budget-constraint formulation. Whereas the original method makes independent per-hypothesis decisions, our framework guarantees adherence to a pre-specified total budget in expectation. Furthermore, we establish that our proposed construction is optimal within a broad and intuitive class of methods, thereby providing stronger theoretical guarantees.
1.2 Our Contributions
To address the aforementioned limitations, we develop a flexible and efficient framework for active hypothesis testing under a computational budget. Our approach leverages inexpensive auxiliary statistics to intelligently allocate computational resources, thereby maximizing statistical power while rigorously respecting budget constraints and maintaining full statistical validity.
Central to the procedure is a control function, guided by an auxiliary statistic, that probabilistically determines whether the true, resource-intensive test statistic should be computed. When the exact statistic is not evaluated, a transformed version of the auxiliary statistic is used in its place, ensuring that a valid $p$-value or $e$-value is produced for every hypothesis. The framework requires only the availability of auxiliary information and a pre-specified budget, making it widely applicable across diverse scientific domains.
Our proposal offers several novel and valuable contributions. First, we establish a formal budget-constrained procedure, guaranteeing that the expected number of expensive computations does not exceed a user-specified limit. This feature makes our framework directly applicable in real-world, resource-constrained settings, providing a formal bridge between statistical objectives and practical limitations. Theoretical guarantees and extensive empirical results confirm that our method achieves tight budget adherence: the actual number of computations closely matches the specified budget, avoiding both wasteful under-utilization and over-expenditure. Second, our framework is model-free, imposing no distributional requirements on the auxiliary statistics. This property is uniquely suited for integrating unstructured information from complex black-box systems, such as LLMs, where the generative process of the auxiliary statistic is unknown or intractable. Finally, we provide rigorous theoretical guarantees for our constructions, showing that our procedure attains optimality for $e$-values and for $p$-values under independence, as well as admissibility for $p$-values under general dependence. This positions our approach as a principled and theoretically sound method, rather than a heuristic.
1.3 Organization
The rest of the paper is organized as follows. Section 2.1 introduces the problem formulation. Section 2.2 presents the active $e$-value framework, and Section 2.3 extends this framework to active $p$-values, providing a dual formulation. Section 2.4 discusses theoretical limitations on the choice of the control function. Section 3 develops the budget-constraint framework and offers practical strategies for selecting the control function based on the system behavior of the auxiliary statistic. Sections 4 and 5 evaluate numerical performance using synthetic data and two real-world case studies: a large-scale GWAS and a clinical application in which auxiliary statistics are generated by an LLM. More discussions and technical proofs are relegated to the supplementary material.
2 A Framework for Active Hypothesis Testing
In many modern scientific applications, such as genomics, drug discovery, or large-scale A/B testing, the number of hypotheses to be tested far exceeds the available computational or experimental resources. This necessitates a principled framework that integrates resource constraints directly into the inferential process. Our goal is to develop a procedure that generates a valid statistical conclusion for every hypothesis while strictly adhering to a pre-specified global budget.
2.1 Problem Formulation
Consider a set of $N$ null hypotheses, $\{H_{0,i}\}_{i=1}^{N}$. For each hypothesis $H_{0,i}$, we have access to two types of statistics:
  1. 1.
    A costly, valid test statistic, denoted generically by $X_{i}$. This represents the “gold-standard” evidence and can be an $e$-value $E_{i}$ or a $p$-value $P_{i}$. The computation or acquisition of $X_{i}$ incurs a significant resource cost.
  2. 2.
    An inexpensive auxiliary statistic, denoted by $X_{i}^{a}$. This statistic (e.g., $E_{i}^{a}$ or $P_{i}^{a}$) is readily available and is assumed to be informative about the exact statistic $X_{i}$, but it may not be statistically valid for formal inference on its own.
Our primary objective is to generate a valid test statistic (an active $e$-value or active $p$-value) for every hypothesis $i\in\{1,\dots,N\}$, while adhering to a pre-specified global budget. We assume that each computation of an expensive statistic $X_{i}$ incurs one unit of cost. The global budget, denoted by $n_{b}$ (where typically $n_{b}\ll N$), represents an upper bound on the expected total number of costly computations.
Formally, let $C_{i}=\mathbb{I}(\text{statistic }X_{i}\text{ is computed})$ be an indicator variable for the decision to compute the expensive statistic for hypothesis $i$. The budget constraint is then given by:
$\mathbb{E}\left[\sum_{i=1}^{N}C_{i}\right]\leq n_{b}.$ (1)
To satisfy this global budget constraint while dynamically allocating resources to the most promising hypotheses, our framework employs hypothesis-specific control functions, $\{h_{i}\}_{i=1}^{N}$. The decision to compute $X_{i}$ is determined by the outcome of a Bernoulli trial with a success probability given by $h_{i}$, which may depend on the full vector of auxiliary statistics $\mathbf{X}^{a}=(X_{1}^{a},\dots,X_{N}^{a})$.
The introduction of these control functions raises the central theoretical question of this work: how can one construct a test statistic that incorporates this probabilistic decision-making while rigorously preserving statistical validity? To answer this, we must first develop the fundamental building block at the level of a single hypothesis. We next define a new object, an “active” statistic, and establish its properties before demonstrating its use in the broader multiple testing setting. We develop this construction in two parallel frameworks, beginning with the active $e$-value.
2.2 Active $e$-value
We begin by considering a single hypothesis. Our goal is to construct an active $e$-value, a composite statistic that leverages an inexpensive auxiliary statistic $E^{a}$ (nonnegative and without further distributional assumptions) to probabilistically decide whether to compute an exact, resource-intensive $e$-value $E$. Recall that a valid $e$-value is any non-negative random variable satisfying $\mathbb{E}_{H_{0}}[E]\leq 1$, where larger values indicate stronger evidence against the null hypothesis.
The decision to compute $E$ is governed by a control function, $h:[0,\infty)\to[0,1]$, which maps the observed value of $E^{a}$ to the probability of computing the exact $e$-value. This probabilistic rule leads to one of two outcomes for the final statistic, as formalized in the following definition.
Definition 1 (Active $e$-value).
The active $e$-value is constructed as:
$E^{\mathrm{active}}=\begin{cases}a(E^{a})&\text{if }U\geq h(E^{a})\\ b(E^{a})\cdot E&\text{if }U<h(E^{a}),\end{cases}$
where $U\sim\text{Uniform}(0,1)$ is independent of $(E^{a},E)$, and $a(\cdot)$ and $b(\cdot)$ are non-negative functions to be designed such that $E^{\mathrm{active}}$ is a valid $e$-value.
Remark 1.
The choice of the multiplicative form $b(E^{a})E$ is deliberate. It is designed to preserve the role of the exact $e$-value, $E$, which is typically a carefully constructed measure of evidence. This structure is intuitive and interpretable, as it simply re-scales the original evidence based on the auxiliary statistic $E^{a}$. We prefer this simple re-scaling to more complex transformations (e.g., $E^{2}$ or other nonlinear functions) that could obscure the relationship between the final statistic and the original $e$-value. Our framework also includes the active $e$-value proposed in Xu et al., 2025b as a special case, with a more detailed comparison provided in Section F of the supplement.
The fundamental theoretical challenge, which we address next, is to determine the conditions on $a(\cdot)$ and $b(\cdot)$ that ensure $E^{\mathrm{active}}$ preserves the $e$-value property, i.e.,
$\mathbb{E}[E^{\mathrm{active}}]=\mathbb{E}[a(E^{a})\cdot(1-h(E^{a}))]+\mathbb{E}[b(E^{a})h(E^{a})\cdot E]\leq 1.$
A natural and intuitive approach to satisfying this inequality is to control each of the two terms in the sum separately. This can be achieved by partitioning the total tolerable expectation of one with a constant $\beta\in[0,1]$, bounding the first term by $\beta$ and the second by $1-\beta$. This decomposition is sufficient, since the two constraints guarantee the total expectation is bounded by $1$. The following theorem provides a complete characterization, showing that this decomposition is not just a convenient strategy but is in fact necessary for the validity of any such active $e$-value construction.
Theorem 1.
For $E^{\mathrm{active}}$ as defined in Definition 1, given the control function $h(\cdot)$, the following are equivalent:
  1. 1.
    $\mathbb{E}[E^{\mathrm{active}}]\leq 1$ for all joint distributions of non-negative random variables $(E^{a},E)$ with $\mathbb{E}[E]\leq 1$.
  2. 2.
    There exists $\beta\in[0,1]$ such that: $\sup_{x\geq 0}a(x)(1-h(x))\leq\beta\text{ and }\sup_{x\geq 0}b(x)h(x)\leq 1-\beta.$
The characterization in Theorem 1 directly informs the optimal design of the functions $a(\cdot)$ and $b(\cdot)$, as demonstrated in the following corollary.
Corollary 1.
For any given $\beta\in[0,1]$ and control function $h(\cdot)$, set:
$a(x)=\frac{\beta}{1-h(x)}\quad\text{and}\quad b(x)=\frac{1-\beta}{h(x)}.$
Then $E^{\mathrm{active}}$ is a valid $e$-value and achieves the tight bound in Theorem 1. In other words, for a fixed $h(\cdot)$ and $\beta$, this construction is optimal in the sense that it is point-wise the largest possible, thus maximizing the resulting $e$-value while preserving validity.
Thus, for a given $\beta$, the active $e$-value construction that is optimal for a fixed control function takes the following explicit form:
$E^{\mathrm{active}}=\begin{cases}\dfrac{\beta}{1-h(E^{a})}&\text{if }U\geq h(E^{a})\\ \dfrac{1-\beta}{h(E^{a})}\cdot E&\text{if }U<h(E^{a}).\end{cases}$ (2)
2.3 Active $p$-value
We now develop an analogous framework for active $p$-values, extending the core principles established for $e$-values. The conceptual setup mirrors the $e$-value case: for a given hypothesis, we have access to an exact, valid $p$-value $P$, and an inexpensive auxiliary statistic $P^{a}$ taking values in $[0,1]$ without further distributional assumptions. We recall that a valid $p$-value $P$ is a random variable satisfying the super-uniformity property under the null hypothesis $H_{0}$: $\mathbb{P}_{H_{0}}(P\leq s)\leq s$ for all $s\in[0,1]$.
Mirroring our approach for $e$-values, the decision to compute the expensive $p$-value is governed by a control function, $h(\cdot):[0,1]\to[0,1]$. This function maps the observed value of the auxiliary statistic $P^{a}$ to the probability of computing the expensive $p$-value, leading to the following definition.
Definition 2 (Active $p$-value).
The active $p$-value, $P^{\mathrm{active}}$, is constructed as follows:
$P^{\mathrm{active}}=\begin{cases}a(P^{a})&\text{if }U\geq h(P^{a})\\ b(P^{a})\cdot P&\text{if }U<h(P^{a}),\end{cases}$
where $U\sim\mathrm{Uniform}(0,1)$ is independent of $(P^{a},P)$, and the functions $a(\cdot)$ and $b(\cdot)$ must be chosen to ensure that $P^{\mathrm{active}}$ is a valid $p$-value.
Remark 2.
Similarly to the $e$-value case, the multiplicative form $b(P^{a})P$ is a deliberate choice. It preserves the original structure of the exact $p$-value $P$, which is often carefully constructed for high power, by simply re-scaling it (e.g., Barber and Ramdas,, 2017; Li and Barber,, 2019; Xia et al.,, 2020; Cai et al.,, 2022). Moreover, our framework encompasses the active $p$-value proposed in Xu et al., 2025b as a special case. Further discussions are provided in Section F of the supplement.
The theoretical challenge is to determine the conditions on $a(\cdot)$ and $b(\cdot)$ that ensure $P^{\mathrm{active}}$ is a valid $p$-value. This requires that for all $s\in[0,1]$,
$\mathbb{P}(P^{\mathrm{active}}\leq s)=\mathbb{E}\left[(1-h(P^{a}))\mathbb{I}\{a(P^{a})\leq s\}+h(P^{a})\mathbb{I}\{b(P^{a})P\leq s\}\right]\leq s.$ (3)
To satisfy this validity condition, we again take a decomposition approach. For a given $\beta\in[0,1]$, we can ensure the total probability is bounded by $s$ if we require that the two terms in the sum are bounded by $\beta s$ and $(1-\beta)s$ respectively:
$\mathbb{E}\left[(1-h(P^{a}))\mathbb{I}\{a(P^{a})\leq s\}\right]$ $\leq\beta s,$ (4)
$\mathbb{E}\left[h(P^{a})\mathbb{I}\{b(P^{a})P\leq s\}\right]$ $\leq(1-\beta)s,$ (5)
for all $s\in[0,1]$. To make the resulting test statistic powerful, we must choose $a(\cdot)$ and $b(\cdot)$ to make the active $p$-value as small as possible. This requires minimizing both of its potential outcomes, $a(P^{a})$ and $b(P^{a})P$, subject to their respective validity constraints.
Remark 3.
The separate constraints (4) and (5) are sufficient, but not necessary, to ensure the active $p$-value satisfies the super-uniformity condition in (3). A counterexample is provided in Section E of the supplement. This contrasts with the $e$-value case in Theorem 1. While our approach imposes a stronger condition than strictly required, it furnishes a tractable framework for constructing a broad class of valid active $p$-values.
We next turn to the optimal form of $a(\cdot)$ from Condition (4). Since $a(P^{a})$ serves as a component of the $p$-value, only values where $a(x)\leq 1$ are meaningful for inference. To maximize statistical power, we seek the point-wise smallest function $a(\cdot)$. The following theorem identifies this optimal choice.
Theorem 2.
Given $\beta$ and $h(\cdot)$, if $a(\cdot)$ satisfies (4) for all distributions of $P^{a}\in[0,1]$ and all $s\in[0,1]$, then $a(x)\geq(1-h(x))/\beta$ whenever $a(x)\leq 1$. Consequently, the choice $a(x)=(1-h(x))/\beta$ is the point-wise smallest selection for the function $a(\cdot)$ under the constraint imposed by (4).
In contrast, the optimal choice of $b(\cdot)$ under Condition (5) is more nuanced, as it is governed by the joint distribution of $P$ and $P^{a}$. For instance, the choice $b(q)=h(q)/(1-\beta)$, which is analogous to the optimal $e$-value construction, fails to satisfy Condition (5) under general dependence (a counterexample is provided in Section D of the supplement). This distinction motivates the need for separate constructions depending on the dependency structure, which we formalize in the following theorem.
Theorem 3.
For fixed $h(\cdot)$ and $\beta$, we have
  1. 1.
    If $P$ and $P^{a}$ are independent, the point-wise smallest $b(\cdot)$ that satisfies (5) is: $b(x)=\frac{h(x)}{1-\beta}.$
  2. 2.
    Under general dependence, an admissible choice for $b(\cdot)$ that satisfies (5) is:
    $b(x)=\frac{\sup_{y}h(y)}{1-\beta}\cdot\mathbb{I}(h(x)>0).$
[…]
Theorems 2 - 3 directly lead to the explicit construction of the active $p$-value. In the following text, the term “active $p$-value” refers to one of these two forms, depending on the dependence between $P$ and $P^{a}$.
Under Independence
When the exact $p$-value $P$ and the auxiliary statistic $P^{a}$ are independent, the active $p$-value takes the following form:
$P^{\mathrm{active}}=\begin{cases}\dfrac{1-h(P^{a})}{\beta}&\text{if }U\geq h(P^{a})\\[10.00002pt] \dfrac{h(P^{a})}{1-\beta}\cdot P&\text{if }U<h(P^{a}).\end{cases}$ (6)
Under General Dependence
To guarantee validity for arbitrary dependence structure between $P$ and $P^{a}$, the construction must adopt a more conservative, uniform scaling factor based on the supremum of the control function. The resulting active $p$-value is:
$P^{\mathrm{active}}=\begin{cases}\dfrac{1-h(P^{a})}{\beta}&\text{if }U\geq h(P^{a})\\[10.00002pt] \dfrac{\sup_{x}h(x)}{1-\beta}\cdot P&\text{if }U<h(P^{a}).\end{cases}$ (7)
A direct comparison of the two forms in (6) and (7) reveals the trade-off between statistical efficiency and robustness. When independence can be assumed, the resulting active $p$-value is smaller (and thus more powerful), as $h(P^{a})\leq\sup_{x}h(x)$. The construction for general dependence pays a price in statistical power to guarantee validity in a wider range of scenarios.
2.4 Admissibility and the Choice of Control Parameters
The active statistic constructions in (2), (6) and (7) depend on the choice of the control function $h(\cdot)$ and the hyperparameter $\beta$. A fundamental theoretical question naturally arises: does a universally optimal configuration exist? That is, can we identify a specific function $h$ and parameter $\beta$ that yield a more powerful test against all alternatives?
To answer this, we formally define statistical domination and admissibility within our framework. Intuitively, one active statistic dominates another if it is always “better”, which means yielding a larger $e$-value or smaller $p$-value regardless of the data realization.
Definition 3 (Domination and Admissibility).
Let $X^{\text{active}}_{h,\beta}$ denote an active statistic (either an active $e$-value or $p$-value) constructed using control function $h$ and parameter $\beta$. We say that $X^{\text{active}}_{h,\beta}$ dominates $X^{\text{active}}_{h^{\prime},\beta^{\prime}}$ if it is strictly more powerful. Formally:
  1. 1.
    For $p$-values: For any valid $p$-value $P$ and auxiliary statistic $P^{a}\in[0,1]$, the inequality $\min\{1,P^{\mathrm{active}}_{h,\beta}\}\leq\min\{1,P^{\mathrm{active}}_{h^{\prime},\beta^{\prime}}\}$ holds almost surely, and strict inequality holds with positive probability for at least one valid input pair.
  2. 2.
    For $e$-values: For any valid $e$-value $E$ and auxiliary statistic $E^{a}\geq 0$, the inequality $E^{\text{active}}_{h,\beta}\geq E^{\text{active}}_{h^{\prime},\beta^{\prime}}$ holds almost surely, and strict inequality holds with positive probability for at least one valid input pair.
An active statistic is admissible if it is not dominated by any other active statistic. We say that a choice of $h(\cdot)$ or $\beta$ is admissible if the resulting active statistic is admissible.
The following propositions establish a key theoretical property of our framework. No single choice of control parameters is universally superior.
Proposition 1 (Admissibility of the Control Function).
Fix $\beta\in(0,1)$. No single control function $h(\cdot)$ uniformly dominates all others. Specifically:
  1. 1.
    For active $e$-values, every choice of $h(\cdot)$ is admissible.
  2. 2.
    For active $p$-values (under both independence and general dependence), every $h(\cdot)$ satisfying $h(\cdot)\geq 1-\beta$ is admissible.
We remark that the constraint $h(\cdot)\geq 1-\beta$ arises because $p$-values greater than 1 are non-informative. Specifically, if $h(x)<1-\beta$, the non-query output $(1-h(x))/\beta$ exceeds 1, providing no evidence against the null.
Proposition 2 (Admissibility of the Hyperparameter).
Assume $h$ is non-trivial (not identically 0 or $1$). No single $\beta\in(0,1)$ uniformly dominates all others. In fact, for any fixed $h(\cdot)$, every active statistic induced by any $\beta\in(0,1)$ is admissible.
The choice of $\beta$ entails a direct trade-off. A larger $\beta$ increases the signal magnitude of the active statistic (yielding a larger $e$-value or smaller $p$-value) when the exact statistic $X$ is not queried, effectively placing more trust in the auxiliary signal. Conversely, a smaller $\beta$ amplifies the result when $X$ is queried. In the absence of specific prior knowledge about the query rate, we recommend $\beta=0.5$ as a robust default, balancing the contribution of the proxy and exact branches.
Finally, while the results in this section focus on a single hypothesis for clarity, they extend naturally to the multivariate setting where $h_{i}$ depends on the full vector of auxiliary statistics $\mathbf{X}^{a}$. We provide the formal extension and proofs of multivariate admissibility in Section B of the supplement.
3 Hypothesis Testing under Budget Constraint
The admissibility results in Section 2.4 establish a fundamental property of our framework: statistical power is not derived from a universally optimal control function, but rather from a data-adaptive strategy that intelligently allocates the global budget $n_{b}$ across the $N$ hypotheses. We now return to the problem formulated in Section 2.1 and present such a strategy in detail.
3.1 A Normalized Allocation Scheme
To connect the global budget $n_{b}$ to the individual decision probabilities $\{h_{i}\}$, we introduce the concept of a utility function, $u_{i}(\cdot)$. For each hypothesis $i$, the utility function $u_{i}:\mathcal{X}^{a}\to\mathbb{R}_{\geq 0}$ maps the auxiliary statistic $X_{i}^{a}$ to a non-negative score that quantifies the “desirability” of computing the exact statistic $X_{i}$. A larger value of $u_{i}(X_{i}^{a})$ indicates a higher priority for allocation of the computational budget.
Given a set of utility functions $\{u_{i}\}_{i=1}^{N}$, we define the control function for each hypothesis via a normalized allocation scheme:
$h_{i}(\mathbf{X}^{a})=\frac{n_{b}\cdot u_{i}(X_{i}^{a})}{\sum_{j=1}^{N}u_{j}(X_{j}^{a})}.$ (8)
This construction is data-driven, allocating a proportionally larger share of the budget to hypotheses with auxiliary statistics deemed more promising. Furthermore, it is designed to satisfy $\sum_{i=1}^{N}h_{i}(\mathbf{X}^{a})=n_{b}$.
Recall that $C_{i}=\mathbb{I}(\text{statistic }X_{i}\text{ is computed})$. Then the expected total cost is bounded by
$\mathbb{E}\Big[\sum_{i=1}^{N}C_{i}\Big]=\sum_{i=1}^{N}\mathbb{E}[\min(1,h_{i}(\mathbf{X}^{a}))]\leq\sum_{i=1}^{N}\mathbb{E}[h_{i}(\mathbf{X}^{a})]=\mathbb{E}\Big[\sum_{i=1}^{N}h_{i}(\mathbf{X}^{a})\Big]=n_{b}.$
So the scheme rigorously adheres to the budget in expectation.
Beyond the expectation-level guarantee, the actual number of computations concentrates sharply around the budget $n_{b}$, as it is a sum of independent Bernoulli random variables. Specifically, by choosing utility functions such that $h_{i}\leq 1$ (cf. Section 3.2), we can use the concentration inequality to show that for any $\delta>0$ and $\epsilon>0$, the total number of expensive computations lies within the interval $n_{b}-\epsilon,n_{b}+\epsilon]$ with probability at least $1-\delta$, provided that the sample size satisfies $N\geq\frac{2\epsilon^{2}}{\log(2/\delta)}$.
3.2 Guidance on Selecting the Utility Functions
Principled strategies for selecting the functional form of $u_{i}$ can lead to substantial gains. The core idea is to encode prior knowledge about the relationship between the auxiliary and exact statistics into the functional form of $u_{i}$.
The Monotonicity Principle.
In most applications, the auxiliary statistic $X_{i}^{a}$ exhibits a consistent, directional relationship with the strength of the evidence against the null. We classify this into two cases:
  1. 1.
    Direct Signal: A signal is considered direct when larger values of $X_{i}^{a}$ are more indicative of the alternative hypothesis. For example, a large $E_{i}^{a}$ may serve as a proxy for a large exact $e$-value $E_{i}$. For direct signals, a non-decreasing utility function $u_{i}(\cdot)$ should be chosen. A natural default choice is the identity function, $u_{i}(x)=x$.
  2. 2.
    Inverse Signal: A signal is inverse when smaller values of $X_{i}^{a}$ are more indicative of the alternative hypothesis (e.g., a small $P_{i}^{a}$ serving as a proxy for an exact $p$-value $P_{i}$). For inverse signals, a non-increasing utility function is appropriate. A standard choice is $u_{i}(x)=1/(x+\epsilon)$, where $\epsilon>0$ is a small constant for numerical stability.
This strategy creates a powerful synergistic effect. Consider the active $e$-value construction (2). Under the alternative, a promising auxiliary statistic (e.g., a large $E_{i}^{a}$ in the direct signal case) will produce a large utility $u_{i}(E_{i}^{a})$, which in turn increases its control value $h_{i}(E_{i}^{a})$. This has two synergistic benefits:
  1. 1.
    It increases the probability of computing the gold-standard $e$-value $E_{i}$, which is also expected to be large.
  2. 2.
    In the event that $E_{i}$ is not computed, the resulting auxiliary-based statistic, $\beta/(1-h_{i}(E_{i}^{a}))$, is also larger, thereby amplifying the evidence from the auxiliary statistic itself.
This dual-benefit mechanism ensures that the budget is efficiently channeled towards maximizing the final evidence against the null.
Practical Refinements.
A crucial practical issue is that the realized number of computations may sometimes fall substantially below the target budget $n_{b}$. To understand why, note that for any given realization of $\mathbf{X}^{a}$, the expected number of computations is $\sum_{i=1}^{N}\min(1,h_{i}(\mathbf{X}^{a}))$, which is always less than or equal to $\sum_{i=1}^{N}h_{i}(\mathbf{X}^{a})=n_{b}$. This inequality becomes strict whenever any control value $h_{i}(\mathbf{X}^{a})>1$. This occurs when the distribution of utility values $\{u_{i}(X_{i}^{a})\}_{i=1}^{N}$ is highly dispersed, with a few outliers being exceptionally large. In such cases, the computational budget cannot be fully utilized.
To address this, we can apply some more range-compressing utility functions to make sure $h_{i}(\mathbf{X}^{a})\leq 1$ for any $i=1,\ldots,N$. A common and effective choice is the logarithm:
$u_{i}(x)=\log(1+x)\quad(\text{for a direct signal}),\quad u_{i}(x)=\log\left(1+\frac{1}{x+\epsilon}\right)\quad(\text{for an inverse signal}).$
This calibration step mitigates the impact of outliers and promotes fuller use of the budget by enabling a more balanced allocation.
3.3 Budgeted Active Inference Algorithm
With the data-adaptive allocation strategy and active statistic constructions in place, we now present the complete, unified algorithm for budgeted active inference. The procedure, detailed in Algorithm 1, takes as input the auxiliary statistics, a global budget, the hyperparameter $\beta$, and user-specified control functions. It returns a valid test statistic for every hypothesis, having rigorously adhered to the budget constraint.
Algorithm 1 A Unified Algorithm for Budgeted Active Inference
0: Auxiliary statistics $\{X_{i}^{a}\}_{i=1}^{N}$, global budget $n_{b}$, hyperparameter $\beta\in(0,1)$, control functions $\{h_{i}(\cdot)\}_{i=1}^{N}$.
1:for each hypothesis $i=1,\dots,N$ do
2:  Draw an independent random variable $U_{i}\sim\mathrm{Uniform}(0,1)$.
3:  if $U_{i}<h_{i}$ then
4:   Compute the exact statistic $X_{i}$.
5:   Construct the active statistic based on $X_{i}$:
  • For $e$-values: $X_{i}^{\mathrm{active}}\leftarrow\frac{1-\beta}{h_{i}}\cdot X_{i}$.
  • For $p$-values: Set $X_{i}^{\mathrm{active}}\leftarrow\min(1,b_{i}\cdot X_{i})$, where the scaling factor $b_{i}$ is:
    • $b_{i}=\frac{h_{i}}{1-\beta}$ (independent),
    • $b_{i}=\frac{\min(1,\sup_{\mathbf{y}\in\mathbb{R}^{N}}h_{i}(\mathbf{y}))}{1-\beta}$ (general).
6:  else
7:   Construct the active statistic without $X_{i}$:
  • For $e$-values: $X_{i}^{\mathrm{active}}\leftarrow\frac{\beta}{1-h_{i}}$.
  • For $p$-values: $X_{i}^{\mathrm{active}}\leftarrow\min(1,\frac{1-h_{i}}{\beta})$.
8:  end if
9:end for
10:Return: The set of active statistics $\{X_{i}^{\mathrm{active}}\}_{i=1}^{N}$.
The set of active statistics $\{X_{i}^{\mathrm{active}}\}_{i=1}^{N}$ produced by Algorithm 1 is designed to be broadly compatible with a wide range of downstream multiple testing procedures, a key advantage of our framework. This design allows researchers to choose the procedure that best suits the form of the statistic produced (whether a $p$-value or an $e$-value), the dependence structure in the data, and the desired power for controlling error metrics such as the False Discovery Rate (FDR, Benjamini and Hochberg, (1995)).
For instance, the resulting active $p$-values can be supplied to a spectrum of methods tailored to different dependency assumptions. These range from the classic Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg,, 1995), which is powerful under independence or PRDS, to the Su procedure, which provides guarantees under the PRDN assumption (Su,, 2018), and the highly robust Benjamini-Yekutieli (BY) procedure (Benjamini and Yekutieli,, 2001) for arbitrary dependence. Moreover, they are compatible with more advanced techniques, including adaptive procedures that estimate the null proportion $\pi_{0}$ to boost power (Storey,, 2002; Storey et al.,, 2004) and sophisticated conditional calibration methods like the dBH procedure (Fithian and Lei,, 2022).
Alternatively, when formulated as $e$-values, our statistics can be integrated with modern $e$-value-based methods, which are particularly appealing for their robustness to complex dependencies. Notable examples include the standard e-BH procedure for arbitrary dependence (Wang and Ramdas,, 2022), enhanced methods that boost power via conditional calibration such as e-BH-CC (Lee and Ren,, 2024), and unifying frameworks like the e-Closure Principle introduced by Xu et al., 2025a , which can offer uniform improvements in power and flexibility. Our framework thus serves as a flexible front-end, compatible with this entire suite of modern statistical machinery.
4 Numerical Experiments
We conduct a series of numerical simulations to evaluate our budgeted active inference framework, comparing its statistical power and efficiency against several baselines under a fixed budget. The utility functions, $\{u_{i}\}_{i=1}^{N}$, are chosen for each experiment as guided by Section 3.2. Unless otherwise specified, we use the identity function $u_{i}(x)=x$ for direct signals and $u_{i}(x)=1/(x+\epsilon)$ with $\epsilon=10^{-8}$ for inverse signals. To handle settings with highly dispersed utilities we may apply the logarithm in the end.
4.1 Competing Methods and Evaluation Metrics
We compare the proposed method “Active-Default” with the following methods:
  1. 1.
    ALL (Oracle). This non-budgeted oracle method computes the exact statistic ($E_{i}$ or $P_{i}$) for all $N$ hypotheses. It serves as an upper bound on statistical power at a fixed cost of $N$ queries.
  2. 2.
    Random. A simple baseline that adheres to the budget by selecting a uniform random subset of $n_{b}$ hypotheses to query. For any hypothesis that is not selected, its active statistic is set to the non-informative value of 1.
  3. 3.
    Xu (Xu et al., 2025b, ). This method makes an independent probabilistic decision for each hypothesis. A Bernoulli trial $T_{i}\sim\mathrm{Bernoulli}(p_{i})$ determines whether to compute the expensive statistic, where the probability $p_{i}$ is a function of the auxiliary statistic and a hyperparameter $\beta$. For $e$-values, the query probability is $p_{i}=\max\{0,1-\beta/E_{i}^{a}\}$, and the final statistic is $E_{i}^{\mathrm{active}}=(1-T_{i})E_{i}^{a}+T_{i}(1-\beta)E_{i}$. For $p$-values, $p_{i}=\max\{0,1-\beta P_{i}^{a}\}$, and the final statistic is $P_{i}^{\mathrm{active}}=(1-T_{i})P_{i}^{a}+T_{i}(1-\beta)^{-1}P_{i}$. Crucially, because these decisions are made independently for each hypothesis, the total number of queries $\sum_{i}T_{i}$ is a random variable and is not constrained by a pre-specified global budget.
  4. 4.
    Active-Xu (Hybrid). An ablation method designed to isolate the benefit of our allocation strategy. It uses the utility function implied by Xu (e.g. $u_{i}(x)=\max(1-\beta/x,0)$ for e-values), but embeds it within our global budget allocation framework.
To evaluate the output statistics, we apply the e-BH procedure to $e$-values and the BY procedure to $p$-values at an FDR level of $\alpha=0.1$, as both are robust to arbitrary dependence structures. Unless otherwise specified, all active methods use the hyperparameter $\beta=0.5$. Our evaluation centers on the trade-off between statistical power and computational cost. We adopt the following standard notation: $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ denote the sets of true null and non-null hypotheses, respectively, with $|\mathcal{H}_{1}|=N_{1}>0$. For a given method, $\mathcal{R}$ is the set of rejected hypotheses.
Statistical Validity and Power.
  • FDR: The expected proportion of false discoveries, defined as $\text{FDR}=\mathbb{E}[V/\max(|\mathcal{R}|,1)]$, where $V=|\mathcal{R}\cap\mathcal{H}_{0}|$. All methods are expected to satisfy $\text{FDR}\leq\alpha$.
  • True Positive Rate (TPR): The expected proportion of true non-nulls correctly rejected, defined as $\text{Power}=\mathbb{E}[S/N_{1}]$, where $S=|\mathcal{R}\cap\mathcal{H}_{1}|$.
Budget-Aware Performance.
Since all methods control FDR, our primary comparison hinges on the efficient use of the computational budget.
  • Queries ($n_{c}$): The total number of expensive computations performed, which directly measures the computational cost and adherence to the budget.
  • Efficiency: The expected number of true discoveries per expensive computation, which captures the return on investment: $\text{Efficiency}=\mathbb{E}\left[{S}/{n_{c}}\right],$ where the ratio is defined as zero if $n_{c}=0$.
4.2 Performance with an Auxiliary Signal
In our first experiment, we assess performance in a scenario where the auxiliary statistic provides a direct but unquantifiable signal about the true effect. We simulate $N=10,000$ hypotheses, each defined by a signal strength parameter $\mu_{i}$. The $i$th null hypothesis is $H_{0,i}:\mu_{i}=0$. The signal strengths $\{\mu_{i}\}_{i=1}^{N}$ are generated independently from a two-component mixture model to create a fraction $\pi$ of non-nulls:
$\mu_{i}\overset{\text{i.i.d.}}{\sim}(1-\pi)\delta_{0}+\pi|\mathcal{N}(0,\tau^{2})|.$
Let $\tau^{2}=2\log N$. From each primary observation $Z_{i}\sim\mathcal{N}(\mu_{i},1)$, we construct a corresponding gold-standard $e$-value and $p$-value: $E_{i}=\exp\left(\lambda Z_{i}-\frac{\lambda^{2}}{2}\right)\text{ and }P_{i}=1-\Phi(Z_{i}),$ where $\lambda=\sqrt{\log(N/\alpha)}$ as recommended in Xu et al., 2025b and $\Phi$ is the standard normal CDF. The corresponding auxiliary statistics, which encode the signal strength $\mu_{i}$, are generated as: $E_{i}^{a}\sim\mathrm{Poisson}(1+\mu_{i})\text{ and }P_{i}^{a}\sim\mathrm{Beta}(1,1+\mu_{i}).$ The Poisson statistic serves as a direct signal for the $e$-value, while the Beta statistic provides an inverse signal for the $p$-value.
We conduct two analyses based on this setup, with a computational budget of $n_{b}=500$. First, to assess performance as a function of signal density, we vary the non-null proportion $\pi$ from 0.05 to 0.3 while holding $\beta=0.5$ fixed. Second, to examine the influence of the hyperparameter $\beta$, we vary it from 0.1 to 0.9 while keeping $\pi=0.1$ fixed.
For our proposed Active-Default method, we use the utility function $u_{i}(x)=x$ for the direct signal ($e$-values) and $u_{i}(x)=\log(1+1/(x+\epsilon))$ for the inverse signal ($p$-values). Given that the statistics are generated independently for each hypothesis, we employ the active $p$-value construction designed for the independent case as in (6). The results for each analysis, averaged over 100 simulations, are presented in Figure 1 and Figure 2, respectively.
Figure 1: Performance comparison as a function of $\pi$ with a budget of $n_{b}=500$. All methods successfully control the FDR at $\alpha=0.1$. Active-Default achieves the highest efficiency.
Figure 2: Performance comparison as a function of the hyperparameter $\beta$ with a budget of $n_{b}=500$. The choice of $\beta$ influences efficiency, but no single value dominates.
The results in Figure 1 clearly demonstrate the practical advantages of our globally budgeted framework. First, the plots confirm that all methods are statistically valid. The FDR panel shows that all procedures maintain the FDR well below the nominal level. The Queries panel confirms that Active-Default, Active-Xu, and Random adhere perfectly to the $n_{b}=500$ budget. In contrast, Xu’s query count grows with $\pi$, exceeding the budget by a factor of 7 to 8 in the $e$-value setting and 4 to 5 in the $p$-value setting.
The central finding lies in the interplay between Power and Efficiency. While the unconstrained Xu and ALL methods achieve higher absolute power, they do so at an enormous computational cost. When performance is measured by efficiency, our proposed Active-Default is the unambiguous winner. Its efficiency grows with $\pi$, indicating that its allocation strategy becomes increasingly effective as the density of true signals increases.
Furthermore, the comparison between Xu and Active-Xu is particularly revealing. By embedding the Xu decision logic within our global budget framework, Active-Xu achieves nearly identical efficiency to its unconstrained counterpart while strictly respecting the budget. This demonstrates the modularity and effectiveness of our allocation scheme. Overall, in a resource-constrained setting where return on investment is paramount, Active-Default provides the best performance.
In Figure 2, we examine the impact of varying the hyperparameter $\beta$ from 0.1 to 0.9 while holding $\pi=0.1$ fixed. We observe that as $\beta$ increases, the efficiency of Active-Default decreases, while the efficiency of Xu and Active-Xu increases. However, as we discussed in Section 2.4, there is no uniformly optimal choice of $\beta$ that dominates across all data-generating mechanisms. The relative performance of different methods depends on the specific characteristics of the problem. Consequently, in practice, we recommend adopting the default choice of $\beta=0.5$, which provides a balanced compromise across a wide range of scenarios.
4.3 Performance with a Noisy Proxy
We next evaluate the methods in a “noisy measurement” setting. This scenario models applications where the auxiliary statistic is not just a simple signal but is itself an “$e$-value” or a “$p$-value” computed from a degraded or noisy version of the primary data. The underlying signal generation remains identical to that in Section 4.2, with hypotheses driven by a signal strength parameter $\mu_{i}$. The key difference lies in the construction of the auxiliary statistic. We create the noisy data $Y_{i}$ by $Y_{i}=Z_{i}+\varepsilon_{i},\text{ where }\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2}).$ From $Z_{i}$ and $Y_{i}$ we construct both $e$-values and $p$-values in parallel:
$E_{i}$ $=\exp\!\left(\lambda Z_{i}-\dfrac{\lambda^{2}}{2}\right)\qquad\text{and}\qquad P_{i}=1-\Phi(Z_{i}),$ (9)
$E_{i}^{a}$ $=\exp\!\left(\lambda Y_{i}-\dfrac{\lambda^{2}}{2}\right)\qquad\text{and}\qquad P_{i}^{a}=1-\Phi(Y_{i}).$ (10)
with $\lambda=\sqrt{\log(N/\alpha)}$. Here $E_{i}^{a}$ is a direct but noisy proxy for $E_{i}$, and $P_{i}^{a}$ is the analogous noisy proxy for $P_{i}$.
We conduct two analyses within this framework. First, we fix the noise standard deviation at a moderate level of $\sigma=1$ and vary the non-null proportion $\pi$ from 0.05 to 0.3. Second, we fix $\pi=0.1$ and vary $\sigma$ from 1 to 5 to assess the methods’ robustness to deteriorating proxy quality. Due to the high dispersion of the auxiliary statistics in this setup, we use the range-compressing utility function $u_{i}(x)=\log(1+x)$ for the $e$-value setting and $\log(1+1/(x+\epsilon))$ for the $p$-value setting to ensure the budget was fully utilized. Here we adopt the active $p$-value constructed for the general dependent case as in (7).
Figure 3: Performance comparison as a function of $\pi$, with a fixed $\sigma=1$.
The results of the first analysis, shown in Figure 3, are highly consistent with our findings from Section 4.2. All methods control the FDR, and our globally budgeted approaches perfectly adhere to the $n_{b}=500$ query limit. The Active-Default method again emerges as the most efficient, with its advantage growing as the proportion of true signals increases.
Figure 4: Performance comparison as a function of $\sigma$, with a fixed $\pi=0.1$.
The second analysis, presented in Figure 4, probes the methods’ robustness. As the noise level $\sigma$ increases, the auxiliary statistic $E_{i}^{a}$ becomes a less reliable indicator of the exact $e$-value $E_{i}$. Consequently, the power and efficiency of Xu, Active-Xu, and Active-Default decline. However, the performance ranking remains stable. Our Active-Default method consistently outperforms the other budget-constrained methods across all noise levels. This demonstrates that even as the quality of the auxiliary information degrades, our framework’s ability to efficiently allocate a fixed budget provides a durable performance advantage.
4.4 Performance with a Correlated Proxy
Our third and final simulation investigates a scenario where the gold-standard and auxiliary statistics share a deeper structural relationship modeled by correlation. This is representative of many real-world problems where a cheap measurement is an indirect but correlated indicator of an expensive one (e.g., gene expression levels and protein abundance).
The underlying signal structure remains the same, while the primary and auxiliary data, $Z_{i}$ and $Y_{i}$, are now drawn from a bivariate normal distribution with correlation $\rho$:
[…]
We then compute $(E_{i},E_{i}^{a})$ and $(P_{i},P_{i}^{a})$ from $(Z_{i},Y_{i})$ via the definitions given in (9) and (10), and the correlation $\rho$ directly controls the quality of both auxiliary channels.
We perform two analyses. First, we fix the correlation at a moderate level, $\rho=0.5$, and vary $\pi$ from 0.05 to 0.3. Second, we fix $\pi=0.1$ and vary $\rho$ from 0.2 to 0.9, assessing how well each method capitalizes on improving proxy quality. As in the default choice, we use $u_{i}(x)=\log(1+x)$ for the $e$-value setting and $u_{i}(x)=\log(1+1/(x+\epsilon))$ for the $p$-value setting. Again, we adopt the active $p$-value constructed for the general dependent case as in (7).
Figure 5: Performance comparison as a function of $\pi$, with a fixed $\rho=0.5$.
The results of the first analysis, displayed in Figure 5, confirm the robustness of our method. The performance patterns are consistent with those observed in the previous, structurally different simulations. Active-Default adheres to the budget while delivering the highest efficiency, with its advantage widening as the proportion of true signals grows.
Figure 6: Performance comparison as a function of $\rho$, with a fixed $\pi=0.1$.
The second analysis, shown in Figure 6, provides a deeper insight into the methods’ behavior. As $\rho$ increases, the auxiliary statistic becomes a more faithful proxy for the gold-standard statistic. This increased information quality allows all active inference methods to improve their power and efficiency. However, Active-Default demonstrates the most significant gains. Its efficiency curve rises more steeply than those of the other methods, highlighting its superior ability to capitalize on high-quality side information. This result shows that our framework not only works well with weak proxies but excels when strong auxiliary data are available, making it an adaptive and powerful tool for budgeted inference.
5 Real-Data Analysis
5.1 Myocardial Infarction GWAS
To demonstrate the practical utility of our framework, we apply it to a common challenge in genomics: leveraging public summary statistics from a GWAS of a related phenotype to guide discovery in a target phenotype under a computational budget. The same framework naturally extends to the same disease across distinct populations or regions, leveraging public GWAS from one group to guide discovery in another (e.g., East Asians vs. Europeans).
Our goal is to identify single-nucleotide polymorphisms (SNPs) associated with myocardial infarction (MI). We use summary statistics from a large GWAS on hypertension (HTN) as inexpensive, auxiliary information. This scenario models a workflow where a research group might repurpose public data to prioritize which SNPs to analyze in their own cohort, thereby saving resources.
We obtained publicly available GWAS summary statistics from the OpenGWAS database. The target phenotype is MI (study ID: ‘ebi-a-GCST90038610’), https://opengwas.io/datasets/ebi-a-GCST90038610 and the auxiliary phenotype is HTN (study ID: ‘ebi-a-GCST90038604’), https://opengwas.io/datasets/ebi-a-GCST90038604.
After aligning the two studies by their SNP identifiers (rsID), we retained $N=9,567,070$ common SNPs. For the $i$th SNP we have its $p$-value from the HTN study, denoted by $P_{i}^{a}$, and its $p$-value from the MI study, denoted by $P_{i}$. A crucial distinction is that $P_{i}^{a}$ is only a valid $p$-value under the null hypothesis of no association with HTN. Under our target null hypothesis (no association with MI), the distribution of $P_{i}^{a}$ is unknown. We therefore treat $\{P_{i}^{a}\}_{i=1}^{N}$ as a set of auxiliary statistics. The MI $p$-values $\{P_{i}\}_{i=1}^{N}$ represent the expensive “gold-standard” evidence whose computation we aim to limit.
We apply our Active-Default framework to this task. Since small $p$-values are the signal of interest (an inverse signal), we use the utility function $u_{i}(x)=\log(1+1/(x+\epsilon))$ with $\epsilon=10^{-8}$. As the two GWAS were conducted on distinct cohorts, we assume the auxiliary and target $p$-values are independent and use the corresponding active $p$-value construction from (6). We include Random, Xu and Active-Xu for comparison.
Since the ground truth is unknown, we establish an oracle set of discoveries to serve as a benchmark, defined as the SNPs rejected by the BY procedure at $\alpha=0.1$ on the full set of $N$ MI $p$-values. We compare the performance of our Active-Default method against Random, Xu, and Active-Xu by measuring their ability to recover these oracle discoveries. For each method, we generate active $p$-values and apply the BY procedure to identify discoveries. Performance is quantified by efficiency, defined as the number of oracle discoveries recovered per MI $p$-value queried. We evaluate this efficiency as the budget, $n_{b}$, is varied as a fraction of the total number of SNPs, with the results summarized in Figure 7.
Figure 7: Performance on the GWAS data analysis. (Left) The number of queried MI $p$-values versus the budget. (Right) The efficiency (oracle discoveries per query) versus the budget.
The left panel of Figure 7 confirms that Active-Default, Active-Xu and the Random baseline precisely adhere to the specified budget, while Xu is unable to provide a budget guarantee. The right panel demonstrates the practical benefit of our approach. At every budget level, the Active-Default method is substantially more efficient than Random and Xu. This indicates that the HTN summary statistics, while not directly valid for inference, provide valuable information for prioritizing the analysis of MI associations, and our framework successfully exploits this information to maximize the return on the computational budget.
To provide external validation, we examine one of the top signals prioritized and discovered by our method, rs1333047. This SNP is located in the 9p21.3 locus, which is one of the most well-established and replicated risk loci for cardiovascular disease. A recent meta-analysis confirmed its strong association with coronary artery disease and MI (Paquette et al.,, 2017), suggesting that testing for this variant could be an important modifier of cardiovascular risk assessment.
5.2 Myocardial Infarction Complications
Our second real-data application also addresses myocardial infarction. We utilize a dataset from Golovenkin et al., (2020) that contains clinical information for 1,700 patients. The primary objective is to predict in-hospital mortality. The original dataset features a multi-class outcome with eight categories: survival and seven distinct causes of death. For our analysis, we simplify this into a testing problem: survival versus mortality, irrespective of the specific cause.
The dataset is partitioned into a training set (800 patients with 672 alive and 128 dead), a calibration set (400 alive patients), and a test set (500 patients with 357 alive and 143 dead). We frame the problem as a multiple hypothesis testing task, where each null hypothesis $H_{0,i}$ posits that patient $i$ will survive.
[…]
$P_{i}=\frac{1+\left|\left\{j\in\mathcal{D}^{\mathrm{cal}}:\hat{s}\left(X_{j}\right)\leq\hat{s}(X_{i})\right\}\right|}{n+1}.$
However, the computation of this exact $p$-value, $P_{i}$, relies on a full feature set, which includes some variables that are costly to acquire. A key example is ‘ZSN_A’, a feature that indicates the presence of chronic heart failure (HF). A definitive diagnosis of HF requires a comprehensive clinical assessment, including symptoms, physical signs, chest X-rays, and echocardiography. The latter, in particular, is an expensive imaging procedure requiring specialized equipment and trained personnel. In our experimental setup, we treat ‘ZSN_A’ as an expensive feature subject to a budget constraint. To perform hypothesis testing under this budget, it is necessary to construct an auxiliary statistic, $P_{i}^{a}$, without access to ‘ZSN_A’.
To generate an auxiliary statistic, $P_{i}^{a}$, without this feature, we leverage an LLM (ChatGPT). We prompt the LLM to impute the ‘ZSN_A’ value for each patient based on their other clinical data. Using this completed dataset, we then compute a proxy conformal $p$-value, $P_{i}^{a}$, to serve as our auxiliary statistic.
Figure 8: Performance on the MI Complications data analysis. All methods control the FDR, and Active-Default achieves the highest efficiency while adhering to the budget.
We then apply our Active-Default framework with a budget of $n_{b}=100$ and a significance level of $\alpha=0.1$ and compare it with Random, Xu, ALL, and Active-Xu baselines. The other settings are the same as in the previous subsection. Discoveries are identified using the BH procedure. While the theoretical validity of BH relies on the PRDS condition, we justify its use here on structural grounds. First, the underlying exact conformal $p$-values satisfy PRDS Bates et al., (2023). Second, although our active $p$-value is not a strictly monotonic transformation of the exact $p$-value, we expect it to remain highly positively correlated with it, thereby preserving the PRDS structure required for FDR control. The results are presented in Figure 8. While all five procedures successfully control the FDR below $\alpha=0.1$, they exhibit marked differences in budget adherence. The budgeted methods (Active-Default, Active-Xu, and Random) precisely respect the $n_{b}=100$ query limit, whereas the unconstrained Xu and ALL methods require significantly more computations. This highlights a clear trade-off between statistical power and computational cost. Although ALL and Xu achieve higher absolute power, our proposed Active-Default demonstrates the highest efficiency, delivering the greatest number of discoveries per query. The value of our global allocation scheme is further underscored by the comparison between Xu and Active-Xu; by enforcing the budget, Active-Xu achieves superior efficiency over its unconstrained counterpart. These findings collectively demonstrate that in resource-constrained settings where efficiency is paramount, our Active-Default framework provides a powerful and principled solution.
6 Discussion
In this work, we developed a general and theoretically grounded framework for active hypothesis testing under a global budget. Our method addresses the challenge of performing statistical inference when the computation of $p$-values or $e$-values is resource-intensive. By using a data-adaptive allocation scheme guided by auxiliary statistics, the framework produces a valid inferential outcome for every hypothesis while ensuring the expected number of expensive computations adheres to a pre-specified limit.
The practical implementation of our framework is guided by the choice of utility functions $\{u_{i}\}$ and the hyperparameter $\beta$. While our admissibility results (Propositions 1 and 2) show that no universally optimal choice exists, we have provided principle-driven guidance that yields effective performance in practice. These considerations also suggest several directions for future work. One direction is the development of data-driven methods for learning better utility functions. One could envision using a held-out calibration dataset to tune the form of $u_{i}(\cdot)$ to maximize a downstream objective, such as the number of discoveries, turning the selection process from a heuristic choice into a formal optimization problem.
Another significant extension would be to adapt our framework from a global budget in expectation to a strict, deterministic budget (i.e., $\sum C_{i}\leq n_{b}$ almost surely). This is a more challenging problem that would likely require techniques from sequential decision-making and online learning, where the decision to query a hypothesis is made adaptively based on the outcomes of previous queries. Finally, our framework could be extended to handle more complex structured inference problems, such as testing hypotheses on a graph or in a sequential, online setting where hypotheses arrive over time.
In conclusion, the budgeted active inference framework presented here offers a flexible method for conducting large-scale hypothesis testing in resource-constrained settings. By formally integrating budget constraints into the inferential process, this work contributes to the development of more efficient and scalable data analysis techniques.