### v2

# Active Hypothesis Testing under Computational Budgets with Applications to GWAS and LLM

Abstract
In large-scale hypothesis testing, computing exact $p$-values or $e$-values is often resource-intensive, creating a need for budget-aware inferential methods. We propose a general framework for active hypothesis testing that leverages inexpensive auxiliary statistics to allocate a global computational budget. For each hypothesis, our data-adaptive procedure probabilistically decides whether to compute the exact test statistic or a transformed proxy, guaranteeing a valid $p$-value or $e$-value while satisfying the exact budget constraint. Theoretical guarantees are established for our constructions, showing that the procedure achieves optimality for $e$-values and for $p$-values under independence, and admissibility for $p$-values under general dependence. Empirical results from simulations and two real-world applications, including a large-scale genome-wide association study (GWAS) and a clinical prediction task leveraging large language models (LLM), demonstrate that our framework improves statistical efficiency under fixed resource limits.
Keywords: Active Learning, Budget-aware Inference, Computational Constraints, $e$-value, Multiple Testing, $p$-value
1 Introduction
The $p$-value and $e$-value (Vovk and Wang, 2021; Ren and Barber, 2023; Ramdas and Wang, 2024) are fundamental tools in statistical inference for quantifying evidence against a null hypothesis. While essential, their exact computation can be prohibitively expensive due to costly experimental procedures or substantial computational demands. This challenge creates a need for inferential methods that operate within a fixed budget. We propose a general framework for active hypothesis testing that addresses this problem directly. Our approach leverages inexpensive and readily available auxiliary statistics, which are derived from cheaper data sources, prior knowledge, or predictive models, to manage a global computational budget. For each hypothesis, a data-adaptive procedure probabilistically decides whether to compute the resource-intensive “gold-standard” statistic. When the exact statistic is not computed, a transformed version of the auxiliary statistic is used in its place, ensuring a valid test statistic for every hypothesis. This framework finds wide applicability across various domains, as illustrated by the following examples.
  1. (a)
    Powerful Prediction Model. Consider a setting where a large, pre-trained prediction model is available to forecast an outcome of interest (Angelopoulos et al., 2023; Motwani and Witten, 2023; Zrnic and Candès, 2024b; Kluger et al., 2025; Ji et al., 2025). The exact $e$-value or $p$-value requires observing the actual outcomes, which may be costly or delayed. A proxy statistic can be rapidly computed by using the model’s predictions in place of the true outcomes. In this setting, the proxy statistic acts as an auxiliary statistic to guide whether observing the true outcomes is worth the cost or delay.
  2. (b)
    Costly vs. Noisy Measurements. In many scientific and industrial domains, a precise measurement is destructive, time-consuming, or financially expensive, while a cheaper but noisier measurement is often available from alternative sensors (Carroll et al., 1995; Fuller, 2009; Grace et al., 2021; Dunbar et al., 2022). For instance, a full genetic assay is costly, but a simple biomarker measurement is not. A valid $p$-value or $e$-value can only be derived from the precise measurement, while the noisy data provides an informative, yet potentially biased, proxy. Our framework formally navigates this trade-off, using the noisy measurement as an informative guide to determine when the budget should be spent on the definitive, precise measurement.
  3. (c)
    Multi-view Learning and Complementary Signals. In many applications, data provide multiple distinct “views” of the same underlying phenomenon, where different perspectives can offer complementary information (Zhang et al., 2011; Sun, 2013; Zhao et al., 2017). A common setting involves two complementary data views. The first view, e.g., genomic or genotype data, is expensive to collect but supports valid computation of $p$-values or $e$-values. The second view, such as routine clinical measurements or gene expression data, is inexpensive to obtain but its null distribution is unknown, making it unsuitable for direct inferential procedures. Despite this limitation, the second view can provide substantial predictive signal. Our framework synthesizes these complementary sources of information by using the cheap data view as a powerful proxy to strategically allocate the budget for computations on the resource-intensive view.
A central challenge addressed in this work is the efficient integration of two types of information: a costly but statistically valid test statistic, and an auxiliary statistic that is inexpensive to obtain but may be less reliable. Formally, we consider a setting with NN hypotheses, each associated with a resource-intensive test statistic that yields valid inference and a cheap auxiliary statistic that may be unreliable. Our objective is to construct a valid test statistic for every hypothesis while ensuring that the number of costly computations strictly adheres to a predetermined global budget.
1.1 Related Work
Prior work on incorporating auxiliary statistics in hypothesis testing has largely focused on improving statistical power, with limited consideration of computational budget constraints. These approaches typically leverage side information to prioritize hypotheses, and are often implemented through weighted multiple testing procedures, which can be interpreted either as re-weighting the $p$-values (Genovese et al., 2006; Ignatiadis et al., 2016; Liu et al., 2016; Barber and Ramdas, 2017; Xia et al., 2020; Cai et al., 2022) or, equivalently, as adaptively adjusting the rejection thresholds (Lei and Fithian, 2018; Zhang et al., 2019; Li and Barber, 2019; Chao and Fithian, 2021; Freestone et al., 2024). Although these approaches improve power, they rely on the assumption that an exact $p$-value is available for every hypothesis and therefore do not tackle the fundamental challenge of high computational or experimental cost. Nevertheless, the weights or prioritization scores generated by some of these approaches can be leveraged as auxiliary statistics within our framework to inform efficient budget allocation. Similarly, two-stage multiple testing procedures use an inexpensive screening stage to filter out unpromising hypotheses (Zehetmayer et al., 2005; Aoshima and Yata, 2011). These methods, however, typically rely on a hard selection rule: hypotheses that fail the initial screening are discarded, and no formal inferential statements are made for them.
A second line of research, which inspires our approach, is active learning, where information is queried selectively to improve efficiency (Cohn et al., 1996; Settles, 2009; Sener and Savarese, 2018; Ren et al., 2021). However, our goal is fundamentally different. The active learning literature, including recent work on acquiring gold-standard labels for statistical inference (Zhang et al., 2021; Zrnic and Candès, 2024a; Cook et al., 2024), has primarily focused on optimizing the collection of labeled data for parameter estimation or model training. While conceptually related, these methods aim to improve the efficiency of data collection, whereas our work focuses on the dynamic decision of whether to compute a test statistic itself.
The work most closely related to ours is the recently proposed proxy computing framework of Xu et al. (2025b), which employs probabilistic queries of exact test statistics to reduce expected computational costs. That approach, however, relies on a fixed test construction and makes query decisions independently for each hypothesis. As a result, it does not provide general optimality guarantees and the total computational cost remains stochastic. We generalize this framework by demonstrating that the construction in Xu et al. (2025b) is a special case of a broader class of valid active statistics. By characterizing this class, we establish the optimality and admissibility theory for active inference and replace the independent query mechanism with a budget-constrained global allocation.
1.2 Our Contributions
To address the aforementioned limitations, we develop a flexible and efficient framework for active hypothesis testing under a computational budget. Our approach leverages inexpensive auxiliary statistics to allocate computational resources in a way that maximizes statistical power, while strictly respecting budget constraints and maintaining statistical validity.
Central to the procedure is a control function, guided by an auxiliary statistic, that probabilistically determines whether the true, resource-intensive test statistic should be computed. When the exact statistic is not evaluated, a transformed version of the auxiliary statistic is used in its place, ensuring that a valid $p$-value or $e$-value is produced for every hypothesis. The framework requires only the availability of auxiliary information and a pre-specified budget, making it widely applicable across diverse scientific domains.
Our work makes several contributions. First, we establish a budget-constrained procedure that guarantees the number of expensive computations exactly matches a user-specified limit on every run. Second, our framework is model-free, imposing no distributional requirements on the auxiliary statistics. This property is uniquely suited for integrating unstructured information from complex black-box systems, such as LLMs, where the generative process of the auxiliary statistic is unknown or intractable. Finally, we provide rigorous theoretical guarantees for our constructions, showing that our procedure attains optimality for $e$-values and for $p$-values under independence, as well as admissibility for $p$-values under general dependence. This positions our approach as a principled and theoretically sound method, rather than a heuristic.
1.3 Organization
The rest of the paper is organized as follows. Section 2.1 introduces the problem formulation. Section 2.2 presents the active $e$-value framework, and Section 2.3 extends this framework to active $p$-values, providing a dual formulation. Section 2.4 discusses theoretical limitations on the choice of the control function. Section 3 develops the budget-constraint framework and offers practical strategies for selecting the control function based on the system behavior of the auxiliary statistic. Sections 4 and 5 evaluate numerical performance using synthetic data and two real-world case studies: a large-scale GWAS and a clinical application in which auxiliary statistics are generated by an LLM. More discussions and technical proofs are relegated to the supplementary material.
2 A Framework for Active Hypothesis Testing
In many modern scientific applications, such as genomics, drug discovery, or large-scale A/B testing, the number of hypotheses to be tested far exceeds the available computational or experimental resources. This necessitates a principled framework that integrates resource constraints directly into the inferential process. Our goal is to develop a procedure that generates a valid statistical conclusion for every hypothesis while strictly adhering to a pre-specified global budget.
2.1 Problem Formulation
Consider a set of NN null hypotheses, $\{H_{0,i}\}_{i=1}^{N}$. For each hypothesis $H_{0,i}$, we have access to two types of statistics:
  1. 1.
    A costly, valid test statistic, denoted generically by $X_{i}$. This represents the “gold-standard” evidence and can be an $e$-value $E_{i}$ or a $p$-value $P_{i}$. The computation or acquisition of $X_{i}$ incurs a significant resource cost.
  2. 2.
    An inexpensive auxiliary statistic, denoted by $X_{i}^{a}$. This statistic (e.g., $E_{i}^{a}$ or $P_{i}^{a}$) is readily available and is assumed to be informative about the exact statistic $X_{i}$, but it may not be statistically valid for formal inference on its own.
Our primary objective is to generate a valid test statistic (an active $e$-value or active $p$-value) for every hypothesis $i\in\{1,\dots,N\}$, while adhering to a pre-specified global budget. We assume that each computation of an expensive statistic $X_{i}$ incurs one unit of cost. The global budget, denoted by $n_{b}$ (where typically $n_{b}\ll N$), represents the total number of costly computations allowed. Formally, let $C_{i}=\mathbb{I}(\text{statistic }X_{i}\text{ is computed})$ be an indicator variable for the decision to compute the expensive statistic for hypothesis ii. The budget constraint is then given by:
$\sum_{i=1}^{N}C_{i}\leq n_{b}.$} (1)
To satisfy this global budget constraint while dynamically allocating resources to the most promising hypotheses, our framework employs hypothesis-specific control functions, $\{h_{i}\}_{i=1}^{N}$. The decision to compute $X_{i}$ is determined by the outcome of a Bernoulli trial with a success probability given by $h_{i}$, which may depend on the full vector of auxiliary statistics $\mathbf{X}^{a}=(X_{1}^{a},\dots,X_{N}^{a})$.
The introduction of these control functions raises the central theoretical question of this work: how can one construct a test statistic that incorporates this probabilistic decision-making while rigorously preserving statistical validity? To answer this, we must first develop the fundamental building block at the level of a single hypothesis. We next define a new object, an “active” statistic, and establish its properties before demonstrating its use in the broader multiple testing setting. We develop this construction in two parallel frameworks, beginning with the active $e$-value.
2.2 Active $e$-value
We begin by considering a single hypothesis. Our goal is to construct an active $e$-value, a composite statistic that leverages an inexpensive auxiliary statistic $E^{a}$ (nonnegative and without further distributional assumptions) to probabilistically decide whether to compute an exact, resource-intensive $e$-value $E$. Recall that a valid $e$-value is any non-negative random variable satisfying $\mathbb{E}_{H_{0}}[E]\leq 1$, where larger values indicate stronger evidence against the null hypothesis.
The decision to compute $E$ is governed by a control function, $h:[0,\infty)\to[0,1]$, which maps the observed value of $E^{a}$ to the probability of computing the exact $e$-value. This probabilistic rule leads to one of two outcomes for the final statistic, as formalized in the following definition.
Definition 1 (Active $e$-value).
The active $e$-value is constructed as:
$E^{\mathrm{active}}=\begin{cases}a(E^{a})&\text{if }U\geq h(E^{a})\\ b(E^{a})\cdot E&\text{if }U<h(E^{a}),\end{cases}$
where $U\sim\text{Uniform}(0,1)$ is independent of $(E^{a},E)$, and $a(\cdot)$ and $b(\cdot)$ are non-negative functions to be designed such that $E^{\mathrm{active}}$ is a valid $e$-value.
Remark 1.
The choice of the multiplicative form $b(E^{a})E$ is deliberate. It is designed to preserve the role of the exact $e$-value, $E$, which is typically a carefully constructed measure of evidence. This structure is intuitive and interpretable, as it simply re-scales the original evidence based on the auxiliary statistic $E^{a}$. We prefer this simple re-scaling to more complex transformations (e.g., $E^{2}$ or other nonlinear functions) that could obscure the relationship between the final statistic and the original $e$-value. Our framework also includes the active $e$-value proposed in Xu et al. (2025b) as a special case, with a more detailed comparison provided in Section G of the supplement.
The fundamental theoretical challenge, which we address next, is to determine the conditions on $a(\cdot)$ and $b(\cdot)$ that ensure $E^{\mathrm{active}}$ preserves the $e$-value property, i.e.,
$\mathbb{E}[E^{\mathrm{active}}]=\mathbb{E}[a(E^{a})\cdot(1-h(E^{a}))]+\mathbb{E}[b(E^{a})h(E^{a})\cdot E]\leq 1.$
A natural and intuitive approach to satisfying this inequality is to control each of the two terms in the sum separately. This can be achieved by partitioning the total tolerable expectation of one with a constant $\beta\in[0,1]$, bounding the first term by $\beta$ and the second by $1-\beta$. This decomposition is sufficient, since the two constraints guarantee the total expectation is bounded by 11. The following theorem provides a complete characterization, showing that this decomposition is not just a convenient strategy but is in fact necessary for the validity of any such active $e$-value construction.
Theorem 1.
For $E^{\mathrm{active}}$ as defined in Definition 1, given the control function $h(\cdot)$, the following two statements are equivalent: (1) $\mathbb{E}[E^{\mathrm{active}}]\leq 1$ for all joint distributions of non-negative random variables $(E^{a},E)$ with $\mathbb{E}[E]\leq 1$; (2) There exists $\beta\in[0,1]$ such that: $\sup_{x\geq 0}a(x)(1-h(x))\leq\beta\text{ and }\sup_{x\geq 0}b(x)h(x)\leq 1-\beta.$
The characterization in Theorem 1 directly informs the optimal design of the functions $a(\cdot)$ and $b(\cdot)$, as demonstrated in the following corollary.
Corollary 1.
For any given $\beta\in[0,1]$ and control function $h(\cdot)$, set:
$a(x)=\frac{\beta}{1-h(x)}\quad\text{and}\quad b(x)=\frac{1-\beta}{h(x)}.$
Then $E^{\mathrm{active}}$ is a valid $e$-value and achieves the tight bound in Theorem 1. In other words, for a fixed $h(\cdot)$ and $\beta$, this construction is optimal in the sense that it is point-wise the largest possible, thus maximizing the resulting $e$-value while preserving validity.
Thus, for a given $\beta$, the active $e$-value construction that is optimal for a fixed control function takes the following explicit form:
$E^{\mathrm{active}}=\begin{cases}\dfrac{\beta}{1-h(E^{a})}&\text{if }U\geq h(E^{a})\\ \dfrac{1-\beta}{h(E^{a})}\cdot E&\text{if }U<h(E^{a}).\end{cases}$ (2)
2.3 Active $p$-value
We now develop an analogous framework for active $p$-values, extending the core principles established for $e$-values. The conceptual setup mirrors the $e$-value case: for a given hypothesis, we have access to an exact, valid $p$-value $P$, and an inexpensive auxiliary statistic $P^{a}$ taking values in $[0,1]$ without further distributional assumptions. We recall that a valid $p$-value $P$ is a random variable satisfying the super-uniformity property under the null hypothesis $H_{0}$: $\mathbb{P}_{H_{0}}(P\leq s)\leq s$ for all $s\in[0,1]$.
Mirroring our approach for $e$-values, the decision to compute the expensive $p$-value is governed by a control function, $h(\cdot):[0,1]\to[0,1]$. This function maps the observed value of the auxiliary statistic $P^{a}$ to the probability of computing the expensive $p$-value, leading to the following definition.
Definition 2 (Active $p$-value).
The active $p$-value, $P^{\mathrm{active}}$, is constructed as follows:
$P^{\mathrm{active}}=\begin{cases}a(P^{a})&\text{if }U\geq h(P^{a})\\ b(P^{a})\cdot P&\text{if }U<h(P^{a}),\end{cases}$
where $U\sim\mathrm{Uniform}(0,1)$ is independent of $(P^{a},P)$, and the functions $a(\cdot)$ and $b(\cdot)$ must be chosen to ensure that $P^{\mathrm{active}}$ is a valid $p$-value.
Remark 2.
Similarly to the $e$-value case, the multiplicative form $b(P^{a})P$ is a deliberate choice. It preserves the original structure of the exact $p$-value $P$, which is often carefully constructed for high power, by simply re-scaling it (e.g., Barber and Ramdas, 2017; Li and Barber, 2019; Xia et al., 2020; Cai et al., 2022). Moreover, our framework encompasses the active $p$-value proposed in Xu et al. (2025b) as a special case. Further discussions are provided in Section G of the supplement.
The theoretical challenge is to determine the conditions on $a(\cdot)$ and $b(\cdot)$ that ensure $P^{\mathrm{active}}$ is a valid $p$-value. This requires that for all $s\in[0,1]$,
$\mathbb{P}(P^{\mathrm{active}}\leq s)=\mathbb{E}\left[(1-h(P^{a}))\mathbb{I}\{a(P^{a})\leq s\}+h(P^{a})\mathbb{I}\{b(P^{a})P\leq s\}\right]\leq s.$ (3)
To satisfy this validity condition, we again take a decomposition approach. For a given $\beta\in[0,1]$, we can ensure the total probability is bounded by $s$ if we require that the two terms in the sum are bounded by $\beta s$ and $(1-\beta)s$ respectively:
$\mathbb{E}\left[(1-h(P^{a}))\mathbb{I}\{a(P^{a})\leq s\}\right]$ $\leq\beta s,$ (4)
$\mathbb{E}\left[h(P^{a})\mathbb{I}\{b(P^{a})P\leq s\}\right]$ $\leq(1-\beta)s,$ (5)
for all $s\in[0,1]$. To make the resulting test statistic powerful, we must choose $a(\cdot)$ and $b(\cdot)$ to make the active $p$-value as small as possible. This requires minimizing both of its potential outcomes, $a(P^{a})$ and $b(P^{a})P$, subject to their respective validity constraints.
Remark 3.
The separate constraints (4) and (5) are sufficient, but not necessary, to ensure the active $p$-value satisfies the super-uniformity condition in (3). A counterexample is provided in Section F of the supplement. This contrasts with the $e$-value case in Theorem 1. While our approach imposes a stronger condition than strictly required, it furnishes a tractable framework for constructing a broad class of valid active $p$-values.
We next turn to the optimal form of $a(\cdot)$ from Condition (4). Since $a(P^{a})$ serves as a component of the $p$-value, only values where $a(x)\leq 1$ are meaningful for inference. To maximize statistical power, we seek the point-wise smallest function $a(\cdot)$. The following theorem identifies this optimal choice.
Theorem 2.
Given $\beta$ and $h(\cdot)$, if $a(\cdot)$ satisfies (4) for all distributions of $P^{a}\in[0,1]$ and all $s\in[0,1]$, then $a(x)\geq(1-h(x))/\beta$ whenever $a(x)\leq 1$. Consequently, the choice $a(x)=(1-h(x))/\beta$ is the point-wise smallest selection for the function $a(\cdot)$ under the constraint imposed by (4).
In contrast, the optimal choice of $b(\cdot)$ under Condition (5) is more nuanced, as it is governed by the joint distribution of $P$ and $P^{a}$. For instance, the choice $b(q)=h(q)/(1-\beta)$, which is analogous to the optimal $e$-value construction, fails to satisfy Condition (5) under general dependence (a counterexample is provided in Section E of the supplement). This distinction motivates the need for separate constructions depending on the dependency structure, which we formalize in the following theorem.
Theorem 3.
For fixed $h(\cdot)$ and $\beta$, we have
  1. 1.
    If $P$ and $P^{a}$ are independent, the point-wise smallest $b(\cdot)$ that satisfies (5) is:
    $b(x)=\frac{h(x)}{1-\beta}.$
  2. 2.
    Under general dependence, an admissible choice for $b(\cdot)$ that satisfies (5) is:
    $b(x)=\frac{\sup_{y}h(y)}{1-\beta}\cdot\mathbb{I}(h(x)>0).$
[…]
Theorems 2 - 3 directly lead to the explicit construction of the active $p$-value. In the following text, the term “active $p$-value” refers to one of these two forms, depending on the dependence between $P$ and $P^{a}$.
Under Independence
When the exact $p$-value $P$ and the auxiliary statistic $P^{a}$ are independent, the active $p$-value takes the following form:
$P^{\mathrm{active}}=\begin{cases}\dfrac{1-h(P^{a})}{\beta}&\text{if }U\geq h(P^{a})\\[10.00002pt] \dfrac{h(P^{a})}{1-\beta}\cdot P&\text{if }U<h(P^{a}).\end{cases}$ (6)
Under General Dependence
To guarantee validity for arbitrary dependence structure between $P$ and $P^{a}$, the construction must adopt a more conservative, uniform scaling factor based on the supremum of the control function. The resulting active $p$-value is:
$P^{\mathrm{active}}=\begin{cases}\dfrac{1-h(P^{a})}{\beta}&\text{if }U\geq h(P^{a})\\[10.00002pt] \dfrac{\sup_{x}h(x)}{1-\beta}\cdot P&\text{if }U<h(P^{a}).\end{cases}$ (7)
A direct comparison of the two forms in (6) and (7) reveals the trade-off between statistical efficiency and robustness. When independence can be assumed, the resulting active $p$-value is smaller (and thus more powerful), as $h(P^{a})\leq\sup_{x}h(x)$. The construction for general dependence pays a price in statistical power to guarantee validity in a wider range of scenarios.
2.4 Admissibility and the Choice of Control Parameters
The active statistic constructions in (2), (6) and (7) depend on the choice of the control function $h(\cdot)$ and the hyperparameter $\beta$. While recent literature (Xu et al., 2025b) has introduced specific functional forms for active statistics, a rigorous theoretical evaluation of whether these or any other choices are optimal has remained absent. This naturally raises a fundamental question: does a universally optimal configuration actually exist? That is, can we identify a specific function $h$ and parameter $\beta$ that yield a strictly more powerful test against all alternatives? To answer this question and address the gap in prior work, we provide a in-depth theoretical investigation into the admissibility of active statistics. We begin by formally defining statistical domination and admissibility within our framework. Intuitively, one active statistic dominates another if it is always “better”, which means yielding a larger $e$-value or smaller $p$-value regardless of the data realization.
Definition 3 (Domination and Admissibility).
Let $X^{\text{active}}_{h,\beta}$ denote an active statistic (either an active $e$-value or $p$-value) constructed using control function $h$ and parameter $\beta$. We say that $X^{\text{active}}_{h,\beta}$ dominates $X^{\text{active}}_{h^{\prime},\beta^{\prime}}$ if it is strictly more powerful. Formally:
  1. 1.
    For $p$-values: For any valid $p$-value $P$ and auxiliary statistic $P^{a}\in[0,1]$, the inequality $\min\{1,P^{\mathrm{active}}_{h,\beta}\}\leq\min\{1,P^{\mathrm{active}}_{h^{\prime},\beta^{\prime}}\}$ holds almost surely, and strict inequality holds with positive probability for at least one valid input pair.
  2. 2.
    For $e$-values: For any valid $e$-value $E$ and auxiliary statistic $E^{a}\geq 0$, the inequality $E^{\text{active}}_{h,\beta}\geq E^{\text{active}}_{h^{\prime},\beta^{\prime}}$ holds almost surely, and strict inequality holds with positive probability for at least one valid input pair.
An active statistic is admissible if it is not dominated by any other active statistic. We say that a choice of $h(\cdot)$ or $\beta$ is admissible if the resulting active statistic is admissible.
The following propositions establish a key theoretical property of our framework. No single choice of control parameters is universally superior.
Proposition 1 (Admissibility of the Control Function).
Fix $\beta\in(0,1)$. No single control function $h(\cdot)$ uniformly dominates all others. Specifically:
  1. 1.
    For active $e$-values, every choice of $h(\cdot)$ is admissible.
  2. 2.
    For active $p$-values (under both independence and general dependence), every $h(\cdot)$ satisfying $h(\cdot)\geq 1-\beta$ is admissible.
We remark that the constraint $h(\cdot)\geq 1-\beta$ arises because $p$-values greater than 1 are non-informative. Specifically, if $h(x)<1-\beta$, the non-query output $(1-h(x))/\beta$ exceeds 1, providing no evidence against the null.
Proposition 2 (Admissibility of the Hyperparameter).
Assume $h$ is non-trivial (not identically 0 or 11). No single $\beta\in(0,1)$ uniformly dominates all others. In fact, for any fixed $h(\cdot)$, every active statistic induced by any $\beta\in(0,1)$ is admissible.
The choice of $\beta$ entails a direct trade-off. A larger $\beta$ increases the signal magnitude of the active statistic (yielding a larger $e$-value or smaller $p$-value) when the exact statistic $X$ is not queried, effectively placing more trust in the auxiliary signal. Conversely, a smaller $\beta$ amplifies the result when $X$ is queried. In the absence of specific prior knowledge about the query rate, we recommend $\beta=0.5$ as a robust default, balancing the contribution of the proxy and exact branches.
Finally, while the results in this section focus on a single hypothesis for clarity, they extend naturally to the multivariate setting where $h_{i}$ depends on the full vector of auxiliary statistics $\mathbf{X}^{a}$. We provide the formal extension and proofs of multivariate admissibility in Section B of the supplement.
3 Hypothesis Testing under Budget Constraint
The admissibility results in Section 2.4 establish a fundamental property of our framework: statistical power is not derived from a universally optimal control function, but rather from a data-adaptive strategy that intelligently allocates the global budget $n_{b}$ across the NN hypotheses. We now return to the problem formulated in Section 2.1 and present such a strategy.
3.1 A Normalized Allocation Scheme
To connect the global budget $n_{b}$ to the individual decision probabilities $\{h_{i}\}$, we introduce the concept of a utility function, $u_{i}(\cdot)$. For each hypothesis $i$, the utility function $u_{i}:\mathcal{X}^{a}\to\mathbb{R}_{\geq 0}$ maps the auxiliary statistic $X_{i}^{a}$ to a non-negative score that quantifies the “desirability” of computing the exact statistic $X_{i}$. A larger value of $u_{i}(X_{i}^{a})$ indicates a higher priority for allocation of the computational budget.
Given a set of utility functions $\{u_{i}\}_{i=1}^{N}$, we define the control function for each hypothesis via a normalized allocation scheme:
$h_{i}(\mathbf{X}^{a})=n_{b}\cdot\frac{u_{i}(X_{i}^{a})}{\sum_{j=1}^{N}u_{j}(X_{j}^{a})}.$ (8)
By construction, this mathematically ensures the exact sum constraint $\sum_{i=1}^{N}h_{i}(\mathbf{X}^{a})=n_{b}$.
3.2 Guidance on Selecting the Utility Functions
Principled strategies for selecting the functional form of $u_{i}$ can lead to substantial gains. The core idea is to encode prior knowledge about the relationship between the auxiliary and exact statistics into the functional form of $u_{i}$.
In most applications, the auxiliary statistic $X_{i}^{a}$ exhibits a consistent, directional relationship with the strength of the evidence against the null. We classify this into two cases:
  1. 1.
    Direct Signal: A signal is considered direct when larger values of $X_{i}^{a}$ are more indicative of the alternative hypothesis. For example, a large $E_{i}^{a}$ may serve as a proxy for a large exact $e$-value $E_{i}$. For direct signals, a non-decreasing utility function $u_{i}(\cdot)$ should be chosen. A natural default choice is the identity function, $u_{i}(x)=x$.
  2. 2.
    Inverse Signal: A signal is inverse when smaller values of $X_{i}^{a}$ are more indicative of the alternative hypothesis (e.g., a small $P_{i}^{a}$ serving as a proxy for an exact $p$-value $P_{i}$). For inverse signals, a non-increasing utility function is appropriate. A standard choice is $u_{i}(x)=1/(x+\epsilon)$, where $\epsilon>0$ is a small constant for numerical stability.
However, if the base utilities are highly skewed, naively computing allocations via the normalized scheme (8) may yield $h_{i}>1$. Simply capping $h_{i}$ at 11 would cause $\sum h_{i}<n_{b}$, and the available resources will not be utilized fully. To guarantee $h_{i}\in[0,1]$, we employ an adaptive transformation applied to the base utilities: $u_{i}(x)=\log(1+(u_{i}^{\text{base}}(x))^{1/k})$, where $k$ is a positive integer. Intuitively, taking the logarithm reduces large differences among the base utilities, and increasing the integer $k$ enforces a progressively stronger compression. Because $n_{b}\leq N$, there always exists an integer $k$ sufficiently large to guarantee $\max_{i}h_{i}\leq 1$. Crucially, this adaptive compression step relies solely on the auxiliary statistics $\{X_{i}^{a}\}$, so it is computationally inexpensive.
This utility selection strategy creates a strong synergy. Consider the active $e$-value construction (2). Under the alternative, a promising auxiliary statistic (e.g., a large $E_{i}^{a}$ in the direct signal case) will produce a large utility $u_{i}(E_{i}^{a})$, which in turn increases its control value $h_{i}(E_{i}^{a})$. This yields two benefits:
  1. 1.
    It increases the probability of computing the gold-standard $e$-value $E_{i}$, which is also expected to be large.
  2. 2.
    In the event that $E_{i}$ is not computed, the resulting auxiliary-based statistic, $\beta/(1-h_{i}(E_{i}^{a}))$, is also larger, thereby amplifying the evidence from the auxiliary statistic itself.
This dual-benefit mechanism ensures that the budget is efficiently channeled towards maximizing the final evidence against the null.
3.3 Budgeted Active Inference Algorithm
Next, a central technical challenge is to ensure that the total number of expensive computations exactly equals $n_{b}$ on every run. Unlike previous methods (Xu et al., 2025b) that rely on independent coin flips which result in random, unpredictable budget utilization, our framework requires a dependent sampling mechanism that correlates the decisions across all NN hypotheses to ensure strict budget adherence. Formally, we seek to sample binary indicators $C_{1},\dots,C_{N}\in\{0,1\}$ conditionally on $\mathbf{X}^{a}$ such that they satisfy two conditions simultaneously: valid marginal selection probabilities, meaning $C_{i}\mid\mathbf{X}^{a}\sim\mathrm{Bernoulli}(h_{i}(\mathbf{X}^{a}))$ for each $i$; and exact global budget adherence, meaning $\sum_{i=1}^{N}C_{i}=n_{b}$. While the theoretical existence of such a joint distribution is guaranteed by Chen et al. (2022), this existence result does not directly yield a practical sampling algorithm. To address this, the next proposition provides an explicit construction of $C_{1},\ldots,C_{N}$ that satisfies these conditions.
Proposition 3.
Suppose $p_{1},\dots,p_{N}\in[0,1]$ satisfy $\sum_{i=1}^{N}p_{i}=n_{b}\in\mathbb{N}$. Let $S_{i}=\sum_{j=1}^{i}p_{j}$ for $i=1,\dots,N$, with $S_{0}=0$ and $U\sim\mathrm{Uniform}(0,1)$. Define
$C_{i}=\lfloor S_{i}-U\rfloor-\lfloor S_{i-1}-U\rfloor,\quad i=1,\dots,N.$ (9)
Then marginally $C_{i}\sim\text{Bernoulli}(p_{i})$ for all $i$, and $\sum_{i=1}^{N}C_{i}=n_{b}$.
We are now ready to present the complete algorithm for budgeted active inference. The procedure, detailed in Algorithm 1, takes as input the auxiliary statistics, a global exact budget, the hyperparameter $\beta$, and user-specified utility functions. It returns a valid test statistic for every hypothesis and rigorously adheres to the constraints.
Algorithm 1 A Unified Algorithm for Budgeted Active Inference
0: Auxiliary statistics $\{X_{i}^{a}\}_{i=1}^{N}$, exact global budget $n_{b}$, hyperparameter $\beta\in(0,1)$, user-specified base utility functions $\{u_{i}(\cdot)\}_{i=1}^{N}$.
1:Step 1: Normalized Allocation
2: Let $u_{i}=u_{i}(X_{i}^{a})$ for $i=1,\dots,N$ and $k=1$.
3: Set $h_{i}=n_{b}\cdot\frac{u_{i}}{\sum_{j}u_{j}}$ for $i=1,\dots,N$.
4:while $\max_{i}h_{i}>1$ do
5:  Adaptively compress utilities: $u_{i}\leftarrow\log(1+(u_{i})^{1/k})$ for $i=1,\dots,N$.
6:  Recompute $h_{i}=n_{b}\cdot\frac{u_{i}}{\sum_{j}u_{j}}$ for $i=1,\dots,N$.
7:  Update $k\leftarrow k+1$.
8:end while
9:Step 2: Exact-Sum Dependent Sampling
10: Compute cumulative limits $S_{i}=\sum_{j=1}^{i}h_{j}$ (with $S_{0}=0$).
11: Sample $U\sim\mathrm{Uniform}(0,1)$.
12:for each hypothesis $i=1,\dots,N$ do
13:  Set indicator boolean $C_{i}=\lfloor S_{i}-U\rfloor-\lfloor S_{i-1}-U\rfloor$.
14:  if $C_{i}=1$ then
15:   Compute the exact primary statistic $X_{i}$.
16:   Construct the active statistic based on $X_{i}$:
  • For $e$-values: $X_{i}^{\mathrm{active}}\leftarrow\frac{1-\beta}{h_{i}}\cdot X_{i}$.
  • For $p$-values: Set $X_{i}^{\mathrm{active}}\leftarrow\min(1,b_{i}\cdot X_{i})$, where the scaling factor $b_{i}$ is:
    • $b_{i}=\frac{h_{i}}{1-\beta}$ (under independence, i.e. $P_{i}\perp\mathbf{P}^{a}$),
    • $b_{i}=\frac{\min(1,\sup_{\mathbf{y}\in\mathbb{R}^{N}}h_{i}(\mathbf{y}))}{1-\beta}$ (under general dependence).
17:  else
18:   Construct the active statistic without $X_{i}$:
  • For $e$-values: $X_{i}^{\mathrm{active}}\leftarrow\frac{\beta}{1-h_{i}}$.
  • For $p$-values: $X_{i}^{\mathrm{active}}\leftarrow\min(1,\frac{1-h_{i}}{\beta})$.
19:  end if
20:end for
21:Return: The set of active statistics $\{X_{i}^{\mathrm{active}}\}_{i=1}^{N}$.
The set of active statistics $\{X_{i}^{\mathrm{active}}\}_{i=1}^{N}$ produced by Algorithm 1 is designed to be broadly compatible with a wide range of downstream multiple testing procedures, a key advantage of our framework. This design allows researchers to choose the procedure that best suits the form of the statistic produced (whether a $p$-value or an $e$-value), the dependence structure in the data, and the desired power for controlling error metrics such as the False Discovery Rate (FDR, Benjamini and Hochberg, 1995).
For instance, the resulting active $p$-values can be supplied to a spectrum of methods tailored to different dependency assumptions. These range from the classic Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg, 1995), which is powerful under independence or PRDS, to the Su procedure, which provides guarantees under the PRDN assumption (Su, 2018), and the highly robust Benjamini-Yekutieli (BY) procedure (Benjamini and Yekutieli, 2001) for arbitrary dependence. Moreover, they are compatible with more advanced techniques, including adaptive procedures that estimate the null proportion $\pi_{0}$ to boost power (Storey, 2002; Storey et al., 2004) and sophisticated conditional calibration methods like the dBH procedure (Fithian and Lei, 2022).
Alternatively, when formulated as $e$-values, our statistics can be integrated with modern $e$-value-based methods, which are particularly appealing for their robustness to complex dependencies. Notable examples include the standard e-BH procedure for arbitrary dependence (Wang and Ramdas, 2022), enhanced methods that boost power via conditional calibration such as e-BH-$C$ (Lee and Ren, 2024), and unifying frameworks like the e-Closure Principle introduced by Xu et al. (2025a), which can offer uniform improvements in power and flexibility. Our framework thus serves as a flexible front-end, compatible with this entire suite of modern statistical machinery.
4 Numerical Experiments
We conduct numerical simulations to evaluate our budgeted active inference framework, comparing its statistical power and efficiency against several baselines under a fixed budget. Across all experiments we use the function $u_{i}(x)=x$ for direct signals and $u_{i}(x)=1/(x+\epsilon)$ for inverse signals as our base utility functions. Any necessary range compression to bound extreme values is automatically handled by the adaptive constraint loop built within Algorithm 1.
4.1 Competing Methods and Evaluation Metrics
We compare Algorithm  1, referred to as “Active-Default”, with the following methods:
  1. 1.
    ALL (Oracle). This non-budgeted oracle method computes the exact statistic ($E_{i}$ or $P_{i}$) for all NN hypotheses. It serves as an upper bound on statistical power at a fixed cost of NN queries.
  2. 2.
    Random. A simple baseline that adheres to the budget by selecting a uniform random subset of $n_{b}$ hypotheses to query. For any hypothesis that is not selected, its active statistic is set to the non-informative value of 1.
  3. 3.
    Xu (Xu et al., 2025b). This method makes an independent probabilistic decision for each hypothesis. A Bernoulli trial $T_{i}\sim\mathrm{Bernoulli}(p_{i})$ determines whether to compute the expensive statistic, where the probability $p_{i}$ is a function of the auxiliary statistic and a hyperparameter $\beta$. For $e$-values, the query probability is $p_{i}=\max\{0,1-\beta/E_{i}^{a}\}$, and the final statistic is $E_{i}^{\mathrm{active}}=(1-T_{i})E_{i}^{a}+T_{i}(1-\beta)E_{i}$. For $p$-values, $p_{i}=\max\{0,1-\beta P_{i}^{a}\}$, and the final statistic is $P_{i}^{\mathrm{active}}=(1-T_{i})P_{i}^{a}+T_{i}(1-\beta)^{-1}P_{i}$. Crucially, because these decisions are made independently for each hypothesis, the total number of queries $\sum_{i}T_{i}$ is a random variable and is not constrained by a pre-specified global budget.
  4. 4.
    Active-Xu (Hybrid). An ablation method designed to isolate the benefit of our allocation strategy. It uses the utility function implied by Xu (e.g. $u_{i}(x)=\max(1-\beta/x,0)$ for e-values), but embeds it within our global budget allocation framework.
To evaluate the output statistics, we apply the e-BH procedure to $e$-values and the BY procedure to $p$-values at an FDR level of $\alpha=0.1$, as both are robust to arbitrary dependence structures. Unless otherwise specified, all active methods use the hyperparameter $\beta=0.5$. Our evaluation centers on the trade-off between statistical power and computational cost. We adopt the following standard notation: $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ denote the sets of true null and non-null hypotheses, respectively, with $|\mathcal{H}_{1}|=N_{1}>0$. For a given method, $\mathcal{R}$ is the set of rejected hypotheses.
Statistical Validity and Power.
  • FDR: The expected proportion of false discoveries, defined as $\text{FDR}=\mathbb{E}[V/\max(|\mathcal{R}|,1)]$, where $V=|\mathcal{R}\cap\mathcal{H}_{0}|$. All methods are expected to satisfy $\text{FDR}\leq\alpha$.
  • True Positive Rate (TPR): The expected proportion of true non-nulls correctly rejected, defined as $\text{Power}=\mathbb{E}[S/N_{1}]$, where $S=|\mathcal{R}\cap\mathcal{H}_{1}|$.
Budget-Aware Performance.
Since all methods control FDR, our primary comparison hinges on the efficient use of the computational budget.
  • Queries ($n_{c}$): The total number of expensive computations performed, which directly measures the computational cost and adherence to the budget.
  • Efficiency: The expected number of true discoveries per expensive computation, which captures the return on investment: $\text{Efficiency}=\mathbb{E}\left[{S}/{n_{c}}\right],$ where the ratio is defined as zero if $n_{c}=0$.
4.2 Performance with an Auxiliary Signal
In this experiment, we assess performance in a scenario where the auxiliary statistic provides a direct but unquantifiable signal about the true effect. Additional simulations are provided in Section C of the Supplement. We simulate $N=10,000$ hypotheses, each defined by a signal strength parameter $\mu_{i}$. The $i$th null hypothesis is $H_{0,i}:\mu_{i}=0$. The signal strengths $\{\mu_{i}\}_{i=1}^{N}$ are generated independently from a two-component mixture model to create a fraction $\pi$ of non-nulls:
$\mu_{i}\overset{\text{i.i.d.}}{\sim}(1-\pi)\delta_{0}+\pi|\mathcal{N}(0,\tau^{2})|.$
Let $\tau^{2}=2\log N$. From each primary observation $Z_{i}\sim\mathcal{N}(\mu_{i},1)$, we construct a corresponding gold-standard $e$-value and $p$-value: $E_{i}=\exp\left(\lambda Z_{i}-\frac{\lambda^{2}}{2}\right)\text{ and }P_{i}=1-\Phi(Z_{i}),$ where $\lambda=\sqrt{\log(N/\alpha)}$ as recommended in Xu et al. (2025b) and $\Phi$ is the standard normal CDF. The corresponding auxiliary statistics, which encode the signal strength $\mu_{i}$, are generated as: $E_{i}^{a}\sim\mathrm{Poisson}(1+\mu_{i})\text{ and }P_{i}^{a}\sim\mathrm{Beta}(1,1+\mu_{i}).$ The Poisson statistic serves as a direct signal for the $e$-value, while the Beta statistic provides an inverse signal for the $p$-value.
We conduct two analyses based on this setup, with a computational budget of $n_{b}=500$. First, to assess performance as a function of signal density, we vary the non-null proportion $\pi$ from 0.05 to 0.3 while holding $\beta=0.5$ fixed. Second, to examine the influence of the $\beta$, we vary it from 0.1 to 0.9 while keeping $\pi=0.1$ fixed. The target FDR level is 0.1. Given that the statistics are generated independently for each hypothesis, we employ the active $p$-value construction designed for the independent case as in (6). The results for each analysis, averaged over 100 simulations, are presented in Figures 1 and 2, respectively.
Figure 1: Performance comparison as a function of $\pi$ with a budget of $n_{b}=500$. All methods successfully control the FDR at $\alpha=0.1$. Active-Default achieves the highest efficiency.
Figure 2: Performance comparison as a function of the hyperparameter $\beta$ with a budget of $n_{b}=500$. The choice of $\beta$ influences efficiency, but no single value dominates.
The results in Figure 1 clearly demonstrate the practical advantages of our globally budgeted framework. First, the plots confirm that all methods are statistically valid. The FDR panel shows that all procedures maintain the FDR well below the nominal level. The Queries panel confirms that Active-Default, Active-Xu, and Random adhere perfectly to the $n_{b}=500$ budget. In contrast, Xu’s query count grows with $\pi$, exceeding the budget by a factor of 7 to 8 in the $e$-value setting and 4 to 5 in the $p$-value setting.
The central finding lies in the interplay between Power and Efficiency. While the unconstrained Xu and ALL methods achieve higher absolute power, they do so at an enormous computational cost. When performance is measured by efficiency, our proposed Active-Default is the unambiguous winner. Its efficiency grows with $\pi$, indicating that its allocation strategy becomes increasingly effective as the density of true signals increases.
Furthermore, the comparison between Xu and Active-Xu is particularly revealing. By embedding the Xu decision logic within our global budget framework, Active-Xu achieves nearly identical efficiency to its unconstrained counterpart while strictly respecting the budget. This demonstrates the modularity and effectiveness of our allocation scheme. Overall, in a resource-constrained setting where return on investment is paramount, Active-Default provides the best performance.
In Figure 2, we examine the impact of varying the hyperparameter $\beta$ from 0.1 to 0.9 while holding $\pi=0.1$ fixed. We observe that as $\beta$ increases, the efficiency of Active-Default decreases, while the efficiency of Xu and Active-Xu increases. However, as we discussed in Section 2.4, there is no uniformly optimal choice of $\beta$ that dominates across all data-generating mechanisms. The relative performance of different methods depends on the specific characteristics of the problem. Consequently, in practice, we recommend adopting the default choice of $\beta=0.5$, which provides a balanced compromise across a wide range of scenarios.
5 Real-Data Analysis
5.1 Myocardial Infarction GWAS
To demonstrate the practical utility of our framework, we apply it to a common challenge in genomics: leveraging public summary statistics from a GWAS of a related phenotype to guide discovery in a target phenotype under a computational budget. The same framework naturally extends to the same disease across distinct populations or regions, leveraging public GWAS from one group to guide discovery in another (e.g., East Asians vs. Europeans).
Our goal is to identify single-nucleotide polymorphisms (SNPs) associated with myocardial infarction (MI). We use summary statistics from a large GWAS on hypertension (HTN) as inexpensive, auxiliary information. This scenario models a workflow where a research group might repurpose public data to prioritize which SNPs to analyze in their own cohort, thereby saving resources.
We obtained publicly available GWAS summary statistics from the OpenGWAS database. The target phenotype is MI (study ID: ‘ebi-a-GCST90038610’), https://opengwas.io/datasets/ebi-a-GCST90038610 and the auxiliary phenotype is HTN (study ID: ‘ebi-a-GCST90038604’), https://opengwas.io/datasets/ebi-a-GCST90038604.
After aligning the two studies by their SNP identifiers (rsID), we retained $N=9,567,070$ common SNPs. For the $i$th SNP we have its $p$-value from the HTN study, denoted by $P_{i}^{a}$, and its $p$-value from the MI study, denoted by $P_{i}$. A crucial distinction is that $P_{i}^{a}$ is only a valid $p$-value under the null hypothesis of no association with HTN. Under our target null hypothesis (no association with MI), the distribution of $P_{i}^{a}$ is unknown. We therefore treat $\{P_{i}^{a}\}_{i=1}^{N}$ as a set of auxiliary statistics. The MI $p$-values $\{P_{i}\}_{i=1}^{N}$ represent the expensive “gold-standard” evidence whose computation we aim to limit.
We apply our Active-Default framework to this task. Since small $p$-values are the signal of interest (an inverse signal), we use the utility function $1/(x+\epsilon)$ with $\epsilon=10^{-8}$ as the base utility function. As the two GWAS were conducted on distinct cohorts, we assume the auxiliary and target $p$-values are independent and use the corresponding active $p$-value construction from (6). We include Random, Xu and Active-Xu for comparison.
Since the ground truth is unknown, we establish an oracle set of discoveries to serve as a benchmark, defined as the SNPs rejected by the BY procedure at $\alpha=0.1$ on the full set of NN MI $p$-values. We compare the performance of our Active-Default method against Random, Xu, and Active-Xu by measuring their ability to recover these oracle discoveries. For each method, we generate active $p$-values and apply the BY procedure to identify discoveries. Performance is quantified by efficiency, defined as the number of oracle discoveries recovered per MI $p$-value queried. We evaluate this efficiency as the budget, $n_{b}$, is varied as a fraction of the total number of SNPs, with the results summarized in Figure 3.
Figure 3: Performance on the GWAS data analysis. (Left) The number of queried MI $p$-values versus the budget. (Right) The efficiency (oracle discoveries per query) versus the budget.
The left panel of Figure 3 confirms that Active-Default and Active-Xu precisely adhere to the specified budget, while Xu is unable to provide a budget guarantee. The right panel demonstrates the practical benefit of our approach. At every budget level, the Active-Default method is substantially more efficient than Random and Xu. This indicates that the HTN summary statistics, while not directly valid for inference, provide valuable information for prioritizing the analysis of MI associations, and our framework successfully exploits this information to maximize the return on the computational budget.
To provide external validation, we examine one of the top signals prioritized and discovered by our method, rs1333047. This SNP is located in the 9p21.3 locus, which is one of the most well-established and replicated risk loci for cardiovascular disease. A recent meta-analysis confirmed its strong association with coronary artery disease and MI (Paquette et al., 2017), suggesting that testing for this variant could be an important modifier of cardiovascular risk assessment.
5.2 Myocardial Infarction Complications
Our second real-data application also addresses myocardial infarction. We utilize a dataset from Golovenkin et al. (2020) that contains clinical information for 1,700 patients. The primary objective is to predict in-hospital mortality. The original dataset features a multi-class outcome with eight categories: survival and seven distinct causes of death. For our analysis, we simplify this into a testing problem: survival versus mortality, irrespective of the specific cause.
The dataset is partitioned into a training set (800 patients with 672 alive and 128 dead), a calibration set (400 alive patients), and a test set (500 patients with 357 alive and 143 dead). We frame the problem as a multiple hypothesis testing task, where each null hypothesis $H_{0,i}$ posits that patient $i$ will survive.
[…]
$P_{i}=\frac{1+\left|\left\{j\in\mathcal{D}^{\mathrm{cal}}:\hat{s}\left(X_{j}\right)\leq\hat{s}(X_{i})\right\}\right|}{n+1}.$
However, the computation of this exact $p$-value, $P_{i}$, relies on a full feature set, which includes some variables that are costly to acquire. A key example is ‘ZSN_A’, a feature that indicates the presence of chronic heart failure (HF). A definitive diagnosis of HF requires a comprehensive clinical assessment, including symptoms, physical signs, chest X-rays, and echocardiography. The latter, in particular, is an expensive imaging procedure requiring specialized equipment and trained personnel. In our experimental setup, we treat ‘ZSN_A’ as an expensive feature subject to a budget constraint. To perform hypothesis testing under this budget, it is necessary to construct an auxiliary statistic, $P_{i}^{a}$, without access to ‘ZSN_A’.
To generate an auxiliary statistic, $P_{i}^{a}$, without this feature, we leverage a large language model, specifically Gemini 3.1 Pro. We prompt the LLM to impute the missing ‘ZSN_A’ value for each patient based on their other clinical data. Using this imputed dataset, we then compute a proxy conformal $p$-value, $P_{i}^{a}$, to serve as our auxiliary statistic. The complete prompt, the full conversation history, and the attached dataset are available at https://gemini.google.com/share/4922ddaad736.
Figure 4: Performance on the MI Complications data analysis. All methods control the FDR, and Active-Default achieves the highest efficiency while adhering to the budget.
We then apply our Active-Default framework with a budget of $n_{b}=100$ and a significance level of $\alpha=0.1$ and compare it with Random, Xu, ALL, and Active-Xu baselines. As with the GWAS analysis, we treat the auxiliary statistic as inverse signals and use the base utility function $1/(x+\epsilon)$. The other settings are the same as in the previous subsection. Discoveries are identified using the BH procedure. While the theoretical validity of BH relies on the PRDS condition, we justify its use here on structural grounds. First, the underlying exact conformal $p$-values satisfy PRDS (Bates et al., 2023). Second, although our active $p$-value is not a strictly monotonic transformation of the exact $p$-value, we expect it to remain highly positively correlated with it, thereby preserving the PRDS structure required for FDR control. The results are presented in Figure 4. While all five procedures successfully control the FDR below $\alpha=0.1$, they exhibit marked differences in budget adherence. The budgeted methods (Active-Default, Active-Xu, and Random) precisely respect the $n_{b}=100$ query limit, whereas the unconstrained Xu and ALL methods require significantly more computations. This highlights a clear trade-off between statistical power and computational cost. Although ALL and Xu achieve higher absolute power, our proposed Active-Default demonstrates the highest efficiency, delivering the greatest number of discoveries per query. The value of our global allocation scheme is further underscored by the comparison between Xu and Active-Xu; by enforcing the budget, Active-Xu achieves superior efficiency over its unconstrained counterpart. These findings collectively demonstrate that in resource-constrained settings where efficiency is paramount, our Active-Default framework provides a powerful and principled solution.
6 Discussion
In this work, we developed a general and theoretically grounded framework for active hypothesis testing under a global budget. Our method addresses the challenge of performing statistical inference when the computation of $p$-values or $e$-values is resource-intensive. By using a data-adaptive allocation scheme guided by auxiliary statistics, the framework produces a valid inferential outcome for every hypothesis while ensuring that the exact number of expensive computations adheres to a pre-specified limit.
The practical implementation of our framework is guided by the choice of utility functions $\{u_{i}\}$ and the hyperparameter $\beta$. While our admissibility results (Propositions 1 and 2) show that no universally optimal choice exists, we have provided guidance that yields effective performance in practice. These considerations also suggest several directions for future work. One direction is the development of data-driven methods for learning better utility functions. One could envision using a held-out calibration dataset to tune the form of $u_{i}(\cdot)$ to maximize a downstream objective, such as the number of discoveries, turning the selection process from a heuristic choice into a formal optimization problem.
Another significant extension would be to handle more complex structured inference problems, such as testing hypotheses on a graph or in a sequential, online setting where hypotheses arrive over time. We discuss some preliminary ideas for the online setting in Supplement H.
In conclusion, the budgeted active inference framework presented here offers a flexible method for conducting large-scale hypothesis testing in resource-constrained settings. By formally integrating budget constraints into the inferential process, this work contributes to the development of more efficient and scalable data analysis techniques.
Declaration of Generative AI
During the preparation of this work, the authors used Gemini 3.1 Pro in order to polish the language and perform professional proofreading to improve the readability of the manuscript. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Supplementary Material to “Active Hypothesis Testing under Computational Budgets with Applications to GWAS and LLM”
This supplement contains the dominance results of direct construction in Section A, admissibility in multivariate setting in Section B, additional numerical experiments in Section C, technical proofs in Section D, relevant counterexamples in Sections E and F, a detailed comparison with the framework of Xu et al. (2025b) in Section G and a discussion about our framework in the online setting in Section H.
Appendix A Dominance of Direct Construction for Active Statistics
In the main text, we present separate constructions for active $p$-values and active $e$-values. This appendix provides a rigorous justification for this approach by demonstrating that these direct constructions are more powerful than indirect methods that rely on converting between different types of statistics. To formalize this comparison, we first establish the mathematical tools used for such conversions.
A.1 The Connection Between $p$-values and $e$-values via Calibrators
To define an indirect construction path (e.g., constructing an active $p$-value from an active $e$-value), we require a principled method for converting between statistic types. This is the role of calibrators.
A $p$-to-$e$ calibrator is a decreasing function $f:[0,\infty)\to[0,\infty]$ that is zero on $(1,\infty)$ such that for any valid $p$-value $P$, the transformed variable $f(P)$ is a valid $e$-value (Vovk and Wang, 2021). Common examples include $f(p)=1/p$ and $f(p)=-\log p$. These calibrators share a fundamental property, formalized in the following lemma.
Lemma 1.
A $p$-to-$e$ calibrator $f$ must satisfy the inequality $f(x)\cdot x\leq 1$ for all $x\in[0,1]$.
This simple bound is the key to proving that indirect, calibrator-based constructions are suboptimal.
Conversely, conversion from $e$-values to $p$-values is more constrained. The standard $e$-to-$p$ calibrator is the reciprocal, $P=1/E$. As shown in Vovk and Wang (2021), this is the only admissible $e$-to-$p$ calibrator, making it the canonical choice for this transformation. Equipped with these definitions, we can now formally compare the direct and indirect construction methods for both active $p$-values and active $e$-values.
A.2 Dominance of the Direct Active $p$-value Construction
We first demonstrate that constructing an active $p$-value directly is strictly more powerful than an indirect approach that converts $p$-values to $e$-values, applies the active $e$-value construction, and then converts back to a $p$-value. Let $P$ be the exact $p$-value and $P^{a}$ be the auxiliary $p$-value. Let $f$ be a $p$-to-$e$ calibrator and $g(x)=1/x$ be the $e$-to-$p$ calibrator.
A.2.1 Indirect Construction (via $e$-values)
The indirect construction of active $p$-value proceeds in three steps:
  1. 1.
    Conversion to $e$-values: Transform the $p$-values to $e$-values: $E=f(P)$ and $E^{a}=f(P^{a})$.
  2. 2.
    Active $e$-value construction: Given a control function $h_{e}$ and a hyperparameter $\beta$, construct the active $e$-value $E_{\mathrm{indirect}}^{\mathrm{active}}$ as defined in the main text (e.g., Equation (2)):
    $E_{\mathrm{indirect}}^{\mathrm{active}}=\begin{cases}\dfrac{\beta}{1-h_{e}(E^{a})}&\text{if }U\geq h_{e}(E^{a})\\ \dfrac{1-\beta}{h_{e}(E^{a})}E&\text{if }U<h_{e}(E^{a}),\end{cases}$
    where $U\sim\mathrm{Uniform}(0,1)$ is independent of all other variables.
  3. 3.
    Conversion back to $p$-value: Invert the resulting active $e$-value to obtain an active $p$-value: $P_{\mathrm{indirect}}^{\mathrm{active}}=g(E_{\mathrm{indirect}}^{\mathrm{active}})=1/E_{\mathrm{indirect}}^{\mathrm{active}}$.
A.2.2 Direct Construction
The direct construction of an active $p$-value, as presented in the main text (e.g., Equation (6) or (7) depending on dependence), yields a statistic $P_{\mathrm{direct}}^{\mathrm{active}}$. For a given control function $h_{p}$ (related to $h_{e}$ via $h_{p}=h_{e}\circ f^{-1}$) and hyperparameter $\beta$, this statistic in independent case is given by:
$P_{\mathrm{direct}}^{\mathrm{active}}=\begin{cases}\dfrac{1-h_{p}(P^{a})}{\beta}&\text{if }U\geq h_{p}(P^{a})\\ \dfrac{h_{p}(P^{a})}{(1-\beta)P}&\text{if }U<h_{p}(P^{a})\end{cases}$
A.2.3 Proof of Dominance
We now show that $P_{\mathrm{direct}}^{\mathrm{active}}\leq P_{\mathrm{indirect}}^{\mathrm{active}}$ under appropriate conditions, implying the direct method yields a more powerful test.
Consider the case where $P$ and $P^{a}$ are independent. The indirect construction yields:
$P_{\mathrm{indirect}}^{\mathrm{active}}=\dfrac{1}{E_{\mathrm{indirect}}^{\mathrm{active}}}=\begin{cases}\dfrac{1-h_{e}(f(P^{a}))}{\beta}&\text{if }U\geq h_{e}(f(P^{a}))\\ \dfrac{h_{e}(f(P^{a}))}{1-\beta}\dfrac{1}{E}&\text{if }U<h_{e}(f(P^{a}))\end{cases}$
Substituting $E=f(P)$ and $h_{p}=h_{e}\circ f^{-1}$, we have $h_{e}(f(P^{a}))=h_{p}(P^{a})$ and $1/E=1/f(P)$.
If $U\geq h_{p}(P^{a})$, then
$P_{\mathrm{indirect}}^{\mathrm{active}}=\dfrac{1-h_{p}(P^{a})}{\beta}.$
In this branch, $P_{\mathrm{direct}}^{\mathrm{active}}=\dfrac{1-h_{p}(P^{a})}{\beta}$, so $P_{\mathrm{direct}}^{\mathrm{active}}=P_{\mathrm{indirect}}^{\mathrm{active}}$.
If $U<h_{p}(P^{a})$, then
$P_{\mathrm{indirect}}^{\mathrm{active}}=\dfrac{h_{p}(P^{a})}{1-\beta}\dfrac{1}{f(P)}.$
For the direct construction in this branch, we have $P_{\mathrm{direct}}^{\mathrm{active}}=\dfrac{h_{p}(P^{a})}{(1-\beta)P}$. To compare, we must relate $P$ and $1/f(P)$. By Lemma 1, we know that $f(P)\cdot P\leq 1$, which implies $P\leq 1/f(P)$. Therefore,
$P_{\mathrm{direct}}^{\mathrm{active}}=\dfrac{h_{p}(P^{a})}{(1-\beta)P}\geq\dfrac{h_{p}(P^{a})}{(1-\beta)(1/f(P))}=\dfrac{h_{p}(P^{a})f(P)}{1-\beta}=P_{\mathrm{indirect}}^{\mathrm{active}}.$
Thus, $P_{\mathrm{direct}}^{\mathrm{active}}\leq P_{\mathrm{indirect}}^{\mathrm{active}}$ in both branches, and strictly so when $U<h_{p}(P^{a})$ and $P<1/f(P)$. This demonstrates that the direct active $p$-value construction is strictly more powerful than the indirect construction under independence.
However, in the general case when there is an arbitrary dependence between $P$ and $P^{a}$, this strict domination relationship no longer holds. The advantage of the direct $p$-value construction above was intrinsically linked to the independence assumption, which permitted a more powerful formulation. Our active $e$-value framework, in contrast, was designed from the outset for robustness under arbitrary dependence.
When the direct $p$-value construction is adapted to handle general dependence, it must adopt a more conservative form, thereby losing the structural advantage it held in the independent setting. At this point, both methods operate under similarly conservative assumptions, so neither holds a fundamental advantage. Their relative performance then depends on the specific data dependence structure, rather than one method being guaranteed to dominate the other.
Conclusion.
The direct construction of active $p$-values yields a statistic that is point-wise no larger than the one obtained via an indirect conversion through $e$-values in the independent case. This implies that the direct method offers strictly greater power, as smaller $p$-values correspond to stronger evidence against the null hypothesis.
A.3 Dominance of the Direct Active $e$-value Construction
We now provide the symmetric argument for active $e$-values, demonstrating that the direct construction is superior to an indirect approach that relies on calibrating through $p$-values. Let $E$ be the exact $e$-value and $E^{a}$ be the auxiliary $e$-value.
A.3.1 Indirect Construction (via $p$-values)
The indirect construction for an active $e$-value proceeds as follows:
  1. 1.
    Conversion to $p$-values: Transform the $e$-values using the reciprocal calibrator: $P=1/E$ and $P^{a}=1/E^{a}$.
  2. 2.
    Active $p$-value construction: Given a control function $h_{p}$ and a hyperparameter $\beta$, construct the active $p$-value $P_{\mathrm{indirect}}^{\mathrm{active}}$ using the formulation from the main text (e.g., Equation (6) or (7)). This yields:
    $P_{\mathrm{indirect}}^{\mathrm{active}}=\begin{cases}\dfrac{1-h_{p}(P^{a})}{\beta}&\text{if }U\geq h_{p}(P^{a})\\ b(P^{a})P&\text{if }U<h_{p}(P^{a})\end{cases}$
    where $b(\cdot)$ is the scaling function defined in Theorem 3 of the main text.
  3. 3.
    Conversion back to $e$-value: Apply a $p$-to-$e$ calibrator $f$ to obtain the final active $e$-value: $E_{\mathrm{indirect}}^{\mathrm{active}}=f(P_{\mathrm{indirect}}^{\mathrm{active}})$.
A.3.2 Direct Construction
The direct construction of an active $e$-value, $E_{\mathrm{direct}}^{\mathrm{active}}$, for a control function $h_{e}$ (related to $h_{p}$ via $h_{e}(x)=h_{p}(1/x)$) is given by:
$E_{\mathrm{direct}}^{\mathrm{active}}=\begin{cases}\dfrac{\beta}{1-h_{e}(E^{a})}&\text{if }U\geq h_{e}(E^{a})\\ \dfrac{1-\beta}{h_{e}(E^{a})}E&\text{if }U<h_{e}(E^{a})\end{cases}$
A.3.3 Proof of Dominance
We now show that $E_{\mathrm{direct}}^{\mathrm{active}}\geq E_{\mathrm{indirect}}^{\mathrm{active}}$, establishing the superior power of the direct method. We analyze the two branches of the construction separately.
If $U\geq h_{e}(E^{a})$, the corresponding indirect $p$-value is $P_{\mathrm{indirect}}^{\mathrm{active}}=\frac{1-h_{p}(P^{a})}{\beta}$. Substituting $P^{a}=1/E^{a}$ and $h_{p}(x)=h_{e}(1/x)$, this becomes $\frac{1-h_{e}(E^{a})}{\beta}$. The final indirect $e$-value is therefore:
$E_{\mathrm{indirect}}^{\mathrm{active}}=f\left(\frac{1-h_{e}(E^{a})}{\beta}\right).$
By Lemma 1, we have $f(x)\leq 1/x$. Applying this gives:
$E_{\mathrm{indirect}}^{\mathrm{active}}\leq\frac{\beta}{1-h_{e}(E^{a})}=E_{\mathrm{direct}}^{\mathrm{active}}.$
Thus, the direct construction dominates in this branch.
If $U<h_{e}(E^{a})$, the indirect $p$-value is $P_{\mathrm{indirect}}^{\mathrm{active}}=b(P^{a})P=b(1/E^{a})/E$. The final indirect $e$-value is:
$E_{\mathrm{indirect}}^{\mathrm{active}}=f\left(\frac{b(1/E^{a})}{E}\right).$
Again applying Lemma 1, we get:
$E_{\mathrm{indirect}}^{\mathrm{active}}\leq\frac{E}{b(1/E^{a})}.$
From Theorem 3 in the main text, any valid choice for $b(\cdot)$ must satisfy $b(P^{a})\geq\frac{h_{p}(P^{a})}{1-\beta}$ (under independence) or a similar lower bound. This implies $\frac{1}{b(P^{a})}\leq\frac{1-\beta}{h_{p}(P^{a})}$. Substituting this into our inequality yields:
$E_{\mathrm{indirect}}^{\mathrm{active}}\leq E\cdot\frac{1-\beta}{h_{p}(1/E^{a})}=E\cdot\frac{1-\beta}{h_{e}(E^{a})}=E_{\mathrm{direct}}^{\mathrm{active}}.$
The direct construction also dominates in this second branch.
Conclusion.
The direct construction of active $e$-values yields a statistic that is point-wise no smaller than the one obtained from an indirect conversion through $p$-values. Since larger $e$-values correspond to stronger evidence against the null, the direct method is provably more powerful and is the preferred approach.
Appendix B Admissibility in Multivariate Setting
In Section 2.4 of the main text, we established the admissibility of active statistics for a single hypothesis, focusing on the scalar control function $h(\cdot)$ and hyperparameter $\beta$. However, under the global budget framework, the decision probabilities for individual hypotheses are coupled through the budget constraint. Consequently, the control function becomes a multivariate vector mapping the full set of auxiliary statistics $\mathbf{X}^{a}=(X^{a}_{1},\dots,X^{a}_{N})$ to a vector of probabilities. This section extends the concept of admissibility to this multivariate, budget-constrained setting. We formalize domination via component-wise vector comparisons and prove that no feasible allocation strategy uniformly dominates another.
To satisfy the global budget constraint, the vector of control functions $\mathbf{h}=(h_{1},\dots,h_{N})$ must satisfy
$\sum_{i=1}^{N}h_{i}(\mathbf{X}^{a})=n_{b}$
for any realization of $\mathbf{X}^{a}$. We denote by $\mathcal{H}$ the set of all valid control function vectors $\mathbf{h}:\mathcal{X}^{N}\to[0,1]^{N}$ that satisfy the above equality. The concept of domination extends naturally to the multivariate case by comparing vectors of active statistics.
Definition B.1 (Multivariate Domination and Admissibility).
Let $\mathbf{X}^{\text{active}}_{\mathbf{h},\boldsymbol{\beta}}$ denote the vector of active statistics induced by a control vector $\mathbf{h}\in\mathcal{H}$ and a hyperparameter vector $\boldsymbol{\beta}\in(0,1)^{N}$.
  1. 1.
    For $p$-values: The vector $\mathbf{X}^{\text{active}}_{\tilde{\mathbf{h}},\tilde{\boldsymbol{\beta}}}$ dominates $\mathbf{X}^{\text{active}}_{\mathbf{h},\boldsymbol{\beta}}$ if, for any valid input, the component-wise inequality
    $\min\{\mathbf{1},\mathbf{X}^{\text{active}}_{\tilde{\mathbf{h}},\tilde{\boldsymbol{\beta}}}\}\leq\min\{\mathbf{1},\mathbf{X}^{\text{active}}_{\mathbf{h},\boldsymbol{\beta}}\}$
    holds almost surely, and there exists at least one valid input distribution such that, with positive probability, the inequality is strict for at least one component.
  2. 2.
    For $e$-values: The vector $\mathbf{X}^{\text{active}}_{\tilde{\mathbf{h}},\tilde{\boldsymbol{\beta}}}$ dominates $\mathbf{X}^{\text{active}}_{\mathbf{h},\boldsymbol{\beta}}$ if, for any valid input, the component-wise inequality
    $\mathbf{X}^{\text{active}}_{\tilde{\mathbf{h}},\tilde{\boldsymbol{\beta}}}\geq\mathbf{X}^{\text{active}}_{\mathbf{h},\boldsymbol{\beta}}$
    holds almost surely, and there exists at least one valid input distribution such that, with positive probability, the inequality is strict for at least one component.
[…]
The following propositions confirm that the phenomena observed in the univariate case persist in the multivariate setting.
Proposition B.1 (Admissibility of the Allocation Strategy).
No single control vector $\mathbf{h}\in\mathcal{H}$ uniformly dominates all others. Specifically, for a fixed hyperparameter vector $\boldsymbol{\beta}\in(0,1)^{N}$, the active statistic vector induced by any $\mathbf{h}\in\mathcal{H}$ is admissible.
Proposition B.2 (Admissibility of the Hyperparameters).
Assume $\mathbf{h}$ is non-degenerate, meaning that for each component $i\in\{1,\dots,N\}$, the function $h_{i}(\cdot)$ is not identically 0 and not identically 11. Then, for any choice of hyperparameters $\boldsymbol{\beta}\in(0,1)^{N}$, the induced active statistic vector $\mathbf{X}^{\text{active}}_{\mathbf{h},\boldsymbol{\beta}}$ is admissible.
Appendix C Additional Numerical Experiments
C.1 Performance with a Correlated Proxy
This simulation investigates a scenario where the gold-standard and auxiliary statistics share a deeper structural relationship modeled by correlation. This is representative of many real-world problems where a cheap measurement is an indirect but correlated indicator of an expensive one (e.g., gene expression levels and protein abundance).
The underlying signal structure remains the same, while the primary and auxiliary data, $Z_{i}$ and $Y_{i}$, are now drawn from a bivariate normal distribution with correlation $\rho$:
[…]
We then compute $(E_{i},E_{i}^{a})$ and $(P_{i},P_{i}^{a})$ from $(Z_{i},Y_{i})$ via the definitions given in (C.1) and (C.2), and the correlation $\rho$ directly controls the quality of both auxiliary channels.
We perform two analyses. First, we fix the correlation at a moderate level, $\rho=0.5$, and vary $\pi$ from 0.05 to 0.3. Second, we fix $\pi=0.1$ and vary $\rho$ from 0.2 to 0.9, assessing how well each method capitalizes on improving proxy quality. Again, we adopt the active $p$-value constructed for the general dependent case as in (7).
Figure C.1: Performance comparison as a function of $\pi$, with a fixed $\rho=0.5$.
The results of the first analysis, displayed in Figure C.1, confirm the robustness of our method. The performance patterns are consistent with those observed in the previous, structurally different simulations. Active-Default adheres to the budget while delivering the highest efficiency, with its advantage widening as the proportion of true signals grows.
Figure C.2: Performance comparison as a function of $\rho$, with a fixed $\pi=0.1$.
The second analysis, shown in Figure C.2, provides a deeper insight into the methods’ behavior. As $\rho$ increases, the auxiliary statistic becomes a more faithful proxy for the gold-standard statistic. This increased information quality allows all active inference methods to improve their power and efficiency. However, Active-Default demonstrates the most significant gains. Its efficiency curve rises more steeply than those of the other methods, highlighting its superior ability to capitalize on high-quality side information. This result shows that our framework not only works well with weak proxies but excels when strong auxiliary data are available, making it an adaptive and powerful tool for budgeted inference.
C.2 Performance with a Noisy Proxy
We next evaluate the methods in a “noisy measurement” setting. This scenario models applications where the auxiliary statistic is not just a simple signal but is itself an “$e$-value” or a “$p$-value” computed from a degraded or noisy version of the primary data. The underlying signal generation remains identical to that in Section 4.2, with hypotheses driven by a signal strength parameter $\mu_{i}$. The key difference lies in the construction of the auxiliary statistic. We create the noisy data $Y_{i}$ by $Y_{i}=Z_{i}+\varepsilon_{i},\text{ where }\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2}).$ From $Z_{i}$ and $Y_{i}$ we construct both $e$-values and $p$-values in parallel:
$E_{i}$ $=\exp\!\left(\lambda Z_{i}-\dfrac{\lambda^{2}}{2}\right)\qquad\text{and}\qquad P_{i}=1-\Phi(Z_{i}),$ (C.1)
$E_{i}^{a}$ $=\exp\!\left(\lambda Y_{i}-\dfrac{\lambda^{2}}{2}\right)\qquad\text{and}\qquad P_{i}^{a}=1-\Phi(Y_{i}).$ (C.2)
with $\lambda=\sqrt{\log(N/\alpha)}$. Here $E_{i}^{a}$ is a direct but noisy proxy for $E_{i}$, and $P_{i}^{a}$ is the analogous noisy proxy for $P_{i}$.
We conduct two analyses within this framework. First, we fix the noise standard deviation at a moderate level of $\sigma=1$ and vary the non-null proportion $\pi$ from 0.05 to 0.3. Second, we fix $\pi=0.1$ and vary $\sigma$ from 1 to 5 to assess the methods’ robustness to deteriorating proxy quality. Here we adopt the active $p$-value constructed for the general dependent case as in (7).
Figure C.3: Performance comparison as a function of $\pi$, with a fixed $\sigma=1$.
The results of the first analysis, shown in Figure C.3, are highly consistent with our findings from Section 4.2. All methods control the FDR, and our globally budgeted approaches perfectly adhere to the $n_{b}=500$ query limit. The Active-Default method again emerges as the most efficient, with its advantage growing as the proportion of true signals increases.
Figure C.4: Performance comparison as a function of $\sigma$, with a fixed $\pi=0.1$.
The second analysis, presented in Figure C.4, probes the methods’ robustness. As the noise level $\sigma$ increases, the auxiliary statistic $E_{i}^{a}$ becomes a less reliable indicator of the exact $e$-value $E_{i}$. Consequently, the power and efficiency of Xu, Active-Xu, and Active-Default decline. However, the performance ranking remains stable. Our Active-Default method consistently outperforms the other budget-constrained methods across all noise levels. This demonstrates that even as the quality of the auxiliary information degrades, our framework’s ability to efficiently allocate a fixed budget provides a durable performance advantage.
Appendix D Technical Proofs
D.1 Proof of Theorem 1
Proof.
The proof proceeds by contradiction. We assume that statement 2 of the theorem is false, meaning no such $\beta\in[0,1]$ exists. This implies that for any $\beta\in[0,1]$, at least one of the two inequalities in statement 2 is violated.
Let us define the quantities AA and $B$ as the suprema of the two components of the expected $e$-value:
$A:=\sup_{x\geq 0}a(x)(1-h(x))\quad\text{and}\quad B:=\sup_{x\geq 0}b(x)h(x).$
Our initial assumption implies that $A+B>1$. To see why, suppose for contradiction that $A+B\leq 1$. We could then choose $\beta=A$. This choice would satisfy both $A\leq\beta$ and $B\leq 1-A=1-\beta$, which contradicts the assumption that no such $\beta$ exists. Thus, it must be that $A+B>1$.
The core of our proof is to construct a specific joint distribution for $(E^{a},E)$ that is valid (i.e., $\mathbb{E}[E]\leq 1$) but for which the active $e$-value construction fails, yielding $\mathbb{E}[E^{\mathrm{active}}]>1$.
Constructing the Counterexample.
Since $A+B>1$, we can fix a small $\delta>0$ such that $A+B>1+2\delta$. By the definition of the supremum, we can find points $x_{1},x_{2}\geq 0$ such that:
$a(x_{1})(1-h(x_{1}))$ $>A-\delta$
$b(x_{2})h(x_{2})$ $>B-\delta$
Let $c_{1}:=a(x_{1})(1-h(x_{1}))$ and $c_{2}:=b(x_{2})h(x_{2})$. From the above, we have $c_{1}+c_{2}>(A-\delta)+(B-\delta)>(1+2\delta)-2\delta=1$.
Now, for any $\epsilon\in(0,1)$, we define the joint distribution of $(E^{a},E)$ as follows:
  • The auxiliary statistic $E^{a}$ is a discrete random variable taking two values:
    $\mathbb{P}(E^{a}=x_{1})=\epsilon\quad\text{and}\quad\mathbb{P}(E^{a}=x_{2})=1-\epsilon.$
  • The exact $e$-value $E$ is conditionally defined based on $E^{a}$:
    $E\mid E^{a}=\begin{cases}0&\text{if }E^{a}=x_{1}\\ \frac{1}{1-\epsilon}&\text{if }E^{a}=x_{2}\end{cases}$
This construction defines a valid joint distribution where the exact $e$-value $E$ has an expectation of 1 under the null, since $\mathbb{E}[E]=\epsilon\cdot 0+(1-\epsilon)\cdot\frac{1}{1-\epsilon}=1$.
Deriving the Contradiction.
We now compute the expectation of the resulting active $e$-value, $E^{\mathrm{active}}$:
$\mathbb{E}[E^{\mathrm{active}}]$ $=\epsilon\cdot\mathbb{E}[E^{\mathrm{active}}\mid E^{a}=x_{1}]+(1-\epsilon)\cdot\mathbb{E}[E^{\mathrm{active}}\mid E^{a}=x_{2}]$
$=\epsilon\left[a(x_{1})(1-h(x_{1}))+b(x_{1})h(x_{1})\cdot 0\right]$
$\quad+(1-\epsilon)\left[a(x_{2})(1-h(x_{2}))+b(x_{2})h(x_{2})\cdot\frac{1}{1-\epsilon}\right]$
$=\epsilon\cdot c_{1}+(1-\epsilon)a(x_{2})(1-h(x_{2}))+c_{2}.$
Since $a(x_{2})(1-h(x_{2}))\geq 0$, we can lower bound this expectation:
$\mathbb{E}[E^{\mathrm{active}}]\geq\epsilon c_{1}+c_{2}.$
Since $c_{1}=a(x_{1})(1-h(x_{1}))\geq 0$ and we have established $c_{1}+c_{2}>1$, there obviously exists some $\epsilon\in(0,1)$ such that $\epsilon c_{1}+c_{2}>1$.
For such an $\epsilon$, we have shown that $\mathbb{E}[E^{\mathrm{active}}]>1$. This contradicts the requirement that $E^{\mathrm{active}}$ must be a valid $e$-value (i.e., $\mathbb{E}[E^{\mathrm{active}}]\leq 1$) for all valid joint distributions of $(E^{a},E)$.
Therefore, our initial assumption must be false, and there must exist a $\beta\in[0,1]$ satisfying the conditions of the theorem. ∎
D.2 Proof of Theorem 2
Proof.
The proof consists of two parts. First, we establish a necessary lower bound for $a(x)$ by considering a specific distribution for $P^{a}$—namely, a point mass at $x$. This forces $a(x)$ to satisfy a point-wise inequality. Second, we verify that the function achieving this lower bound is indeed sufficient to satisfy the validity condition for any general distribution of $P^{a}$.
We first show the necessity. Fix any $x\in[0,1]$ such that $a(x)\leq 1$. Consider a point-mass distribution for the auxiliary statistic, $P^{a}\equiv x$. In this case, Condition (4) must hold for all $s\in[0,1]$. Specifically, choosing $s=a(x)$ (which is valid since $a(x)\leq 1$), the condition becomes:
$\mathbb{E}\left[(1-h(P^{a}))\mathbb{I}\{a(P^{a})\leq a(x)\}\right]=(1-h(x))\cdot 1\leq\beta a(x).$
This inequality implies $a(x)\geq(1-h(x))/\beta$.
Next, we show the sufficiency of the choice $a(x)=(1-h(x))/\beta$. Substituting this form into the left-hand side of (4), we have:
$\mathbb{E}\left[(1-h(P^{a}))\mathbb{I}\left\{\frac{1-h(P^{a})}{\beta}\leq s\right\}\right]=\mathbb{E}\left[(1-h(P^{a}))\mathbb{I}\{1-h(P^{a})\leq\beta s\}\right]\leq\beta s.$
D.3 Proof of Theorem 3
Proof.
We prove the two parts of the theorem separately. First, we establish the point-wise optimal choice for $b(\cdot)$ under the assumption of independence. Second, we provide and verify an admissible choice for $b(\cdot)$ for the general case of arbitrary dependence.
Part 1: point-wise Optimal Choice under Independence.
We begin by establishing a necessary lower bound that any valid function $b(\cdot)$ must satisfy. Consider a fixed auxiliary value $P^{a}=q$ and an independent exact $p$-value $P\sim\mathrm{Uniform}(0,1)$. For condition (5) to hold, we must have:
$\mathbb{E}\left[h(q)\mathbb{I}\{b(q)P\leq s\}\right]=h(q)\cdot\mathbb{P}\left(P\leq\frac{s}{b(q)}\right)=h(q)\cdot\min\left\{1,\frac{s}{b(q)}\right\}\leq(1-\beta)s.$
For any $s$ small enough such that $s/b(q)\leq 1$, this inequality simplifies to $h(q)\cdot\frac{s}{b(q)}\leq(1-\beta)s$. This directly implies the necessary condition:
$b(q)\geq\frac{h(q)}{1-\beta}.$
Next, we verify that the function $b(q)=\frac{h(q)}{1-\beta}$ satisfies condition (5) when $P\perp P^{a}$. The expectation becomes $\mathbb{E}[h(P^{a})\cdot\min\{1,\frac{(1-\beta)s}{h(P^{a})}\}]=\mathbb{E}[\min\{h(P^{a}),(1-\beta)s\}]\leq(1-\beta)s$. Since $b(q)=\frac{h(q)}{1-\beta}$ meets the necessary lower bound, it is the point-wise smallest valid choice, and thus optimal under the independence assumption. As shown in Section E, this choice is not valid under general dependence.
Part 2: Admissible Choice under General Dependence.
For the general case, we propose the choice $b^{*}(q):=\frac{\sup_{x}h(x)}{1-\beta}\mathbb{I}\{h(q)>0\}$. We prove its suitability by establishing its validity and then its admissibility.
Validity.
Let $M:=\sup_{x}h(x)$. We must show that $\mathbb{E}[h(P^{a})\mathbb{I}\{b^{*}(P^{a})P\leq s\}]\leq(1-\beta)s$.
$\mathbb{E}[h(P^{a})\mathbb{I}\{b^{*}(P^{a})P\leq s\}]$ $=\mathbb{E}\left[h(P^{a})\mathbb{I}\left\{\frac{M}{1-\beta}P\leq s\right\}\right]$
$\leq\mathbb{E}\left[M\cdot\mathbb{I}\left\{\frac{M}{1-\beta}P\leq s\right\}\right]$ (since $h(P^{a})\leq M$)
$=M\cdot\mathbb{P}\left(P\leq\frac{(1-\beta)s}{M}\right)$
$\leq M\cdot\frac{(1-\beta)s}{M}$ (since $P$ is super-uniform)
$=(1-\beta)s.$
Thus, the choice $b^{*}(q)$ is valid for any joint distribution of $(P,P^{a})$.
Admissibility.
We prove admissibility by contradiction. Assume $b^{*}(q)$ is not admissible. Then there must exist another valid function, $\tilde{b}(q)$, that dominates $b^{*}(q)$. This means:
  1. 1.
    $\tilde{b}(q)\leq b^{*}(q)$ for all $q\in[0,1]$.
  2. 2.
    There exists at least one point $q_{0}$ where $\tilde{b}(q_{0})<b^{*}(q_{0})$.
The second condition implies $h(q_{0})>0$ (otherwise $b^{*}(q_{0})=0$, contradicting the non-negativity of $\tilde{b}$). Since $b^{*}(q_{0})=M/(1-\beta)$, we can express the strict inequality as $\tilde{b}(q_{0})=\frac{M-\delta}{1-\beta}$ for some $\delta>0$.
By the definition of the supremum, for any $\epsilon>0$, there exists a point $q_{1}$ such that $h(q_{1})>M-\epsilon$. We construct a joint distribution for $(P,P^{a})$ parameterized by a constant $p\in(0,1)$ to be chosen later:
  • Let $\mathbb{P}(P^{a}=q_{1})=p$ and $\mathbb{P}(P^{a}=q_{0})=1-p$.
  • Let the conditional distribution of $P$ be $P\mid(P^{a}=q_{1})\sim\mathrm{Uniform}(0,p)$ and $P\mid(P^{a}=q_{0})\sim\mathrm{Uniform}(p,1)$. This ensures the marginal distribution of $P$ is exactly $\mathrm{Uniform}(0,1)$.
We analyze the validity constraint for $\tilde{b}(q)$ under this specific distribution:
$\mathbb{E}[h(P^{a})\mathbb{I}\{\tilde{b}(P^{a})P\leq s\}]$ $=p\cdot h(q_{1})\mathbb{P}\left(\mathrm{Unif}(0,p)\leq\frac{s}{\tilde{b}(q_{1})}\right)+(1-p)h(q_{0})\mathbb{P}\left(\mathrm{Unif}(p,1)\leq\frac{s}{\tilde{b}(q_{0})}\right).$
We strategically set $p:=\frac{(1-\beta)s}{M}$. From the dominance assumption, $\tilde{b}(q_{1})\leq b^{*}(q_{1})\leq M/(1-\beta)$, implying $s/\tilde{b}(q_{1})\geq s(1-\beta)/M=p$. Thus, the first probability evaluates exactly to 11. Using $h(q_{1})>M-\epsilon$, the expectation becomes:
$\mathbb{E}[h(P^{a})\mathbb{I}\{\tilde{b}(P^{a})P\leq s\}]>p(M-\epsilon)+(1-p)h(q_{0})\mathbb{P}\left(\mathrm{Unif}(p,1)\leq\frac{s}{\tilde{b}(q_{0})}\right).$
For the second term, we evaluate the upper bound inside the probability using $\tilde{b}(q_{0})=\frac{M-\delta}{1-\beta}$:
$\frac{s}{\tilde{b}(q_{0})}=\frac{s(1-\beta)}{M-\delta}>\frac{s(1-\beta)}{M}=p.$
Since $s/\tilde{b}(q_{0})>p$, the probability is strictly positive. Specifically, for sufficiently small $s$ such that $s/\tilde{b}(q_{0})\leq 1$, we have:
$\mathbb{P}\left(\mathrm{Unif}(p,1)\leq\frac{s}{\tilde{b}(q_{0})}\right)=\frac{1}{1-p}\left(\frac{s(1-\beta)}{M-\delta}-p\right)=\frac{p}{1-p}\left(\frac{M}{M-\delta}-1\right)=\frac{p}{1-p}\cdot\frac{\delta}{M-\delta}.$
Substituting $pM=(1-\beta)s$ and the evaluated probability back into the expectation:
$\mathbb{E}[h(P^{a})\mathbb{I}\{\tilde{b}(P^{a})P\leq s\}]$ $>\frac{(1-\beta)s}{M}(M-\epsilon)+(1-p)h(q_{0})\frac{p}{1-p}\frac{\delta}{M-\delta}$
$=(1-\beta)s-\epsilon\frac{(1-\beta)s}{M}+\underbrace{h(q_{0})\frac{(1-\beta)s}{M}\frac{\delta}{M-\delta}}_{:=\Delta}.$
Notice that $\Delta$ depends solely on $M,\delta,\beta,s,$ and $h(q_{0})$ and we can select an $\epsilon$ such that $0<\epsilon<\Delta\cdot\frac{M}{(1-\beta)s}$. With this choice of $\epsilon$, we have:
$\mathbb{E}[h(P^{a})\mathbb{I}\{\tilde{b}(P^{a})P\leq s\}]>(1-\beta)s.$
This strictly violates the validity requirement for $\tilde{b}(q)$. Therefore, no such dominating function $\tilde{b}(q)$ can exist, establishing that $b^{*}(q)$ is admissible. ∎
D.4 Proof of Proposition 1
Proof.
We prove the two parts of the proposition—the admissibility of the control function $h(\cdot)$ for the $e$-value setting and the $p$-value setting. Our proof strategy is to construct specific distributions and events where any two distinct choices outperform each other, thereby demonstrating that no choice can be dominated.
Part 1: $e$-value setting.
Let $h_{1}$ and $h_{2}$ be two distinct control functions. Since they are distinct, there must exist a point $x_{0}\geq 0$ where their values differ. Without loss of generality, assume $h_{1}(x_{0})>h_{2}(x_{0})$.
To show that neither function can dominate the other, we analyze a simple setting where the auxiliary statistic is fixed: $\mathbb{P}(E^{a}=x_{0})=1$. Let $E_{1}^{\text{active}}$ and $E_{2}^{\text{active}}$ be the active $e$-values generated using $h_{1}$ and $h_{2}$, respectively. The outcome depends on the draw of $U\sim\mathrm{Uniform}(0,1)$.
First, consider the event $U\in[h_{1}(x_{0}),1)$, which occurs with positive probability if $h_{1}(x_{0})<1$. On this event, we have $U\geq h_{1}(x_{0})>h_{2}(x_{0})$, so the proxy-based branch is chosen for both constructions. The resulting $e$-values are $E_{1}^{\text{active}}=\frac{\beta}{1-h_{1}(x_{0})}$ and $E_{2}^{\text{active}}=\frac{\beta}{1-h_{2}(x_{0})}$. Since $h_{1}(x_{0})>h_{2}(x_{0})$, it follows that $1-h_{1}(x_{0})<1-h_{2}(x_{0})$, which implies $E_{1}^{\text{active}}>E_{2}^{\text{active}}$.
Second, consider the event $U\in[0,h_{2}(x_{0}))$, which occurs with positive probability if $h_{2}(x_{0})>0$. On this event, $U<h_{2}(x_{0})<h_{1}(x_{0})$, so the exact $e$-value branch is chosen for both. The $e$-values are $E_{1}^{\text{active}}=\frac{1-\beta}{h_{1}(x_{0})}E$ and $E_{2}^{\text{active}}=\frac{1-\beta}{h_{2}(x_{0})}E$. Given that $h_{1}(x_{0})>h_{2}(x_{0})$, we have $\frac{1}{h_{1}(x_{0})}<\frac{1}{h_{2}(x_{0})}$, and thus $E_{1}^{\text{active}}<E_{2}^{\text{active}}$ for any $E$ with positive mass on $(0,\infty)$.
Since we have identified mutually exclusive events with positive probability where each function produces a strictly larger $e$-value, neither can uniformly dominate the other (excluding trivial boundary cases). Therefore, every choice of $h(\cdot)$ is admissible.
Part 2: $p$-value setting.
Let $h_{1}$ and $h_{2}$ be two distinct control functions, and assume without loss of generality that for some point $x_{0}$ we have $h_{1}(x_{0})>h_{2}(x_{0})$. Moreover, assume $h_{1}$ and $h_{2}$ are greater than $1-\beta$. To prove admissibility, we show that neither function can dominate the other by constructing scenarios where each produces a strictly smaller (i.e., better) active $p$-value.
Consider a simple setting where the auxiliary statistic is fixed, $\mathbb{P}(P^{a}=x_{0})=1$. The outcome depends on the draw of $U\sim\mathrm{Uniform}(0,1)$.
First, consider the event $U\in[h_{2}(x_{0}),h_{1}(x_{0}))$. In this case, the resulting active $p$-values are
$P_{1}^{\text{active}}=\frac{h_{1}(x_{0})}{1-\beta}P\quad\text{and}\quad P_{2}^{\text{active}}=\frac{1-h_{2}(x_{0})}{\beta}.$
Taking $P\sim\mathrm{Uniform}(0,1)$ and noting that $h_{1}(x_{0})>h_{2}(x_{0})\geq 1-\beta$, we have $\frac{1-\beta}{\beta}\cdot\frac{1-h_{2}(x_{0})}{h_{1}(x_{0})}\in(0,1)$. Consequently, with probability $1-\frac{1-\beta}{\beta}\cdot\frac{1-h_{2}(x_{0})}{h_{1}(x_{0})}$, we have
$P>\frac{1-\beta}{\beta}\cdot\frac{1-h_{2}(x_{0})}{h_{1}(x_{0})},$
which is equivalent to $\frac{h_{1}(x_{0})}{1-\beta}P>\frac{1-h_{2}(x_{0})}{\beta}$, i.e., $P_{1}^{\text{active}}>P_{2}^{\text{active}}$. Similarly, with probability $\frac{1-\beta}{\beta}\cdot\frac{1-h_{2}(x_{0})}{h_{1}(x_{0})}$, we have
$P<\frac{1-\beta}{\beta}\cdot\frac{1-h_{2}(x_{0})}{h_{1}(x_{0})},$
which is equivalent to $\frac{h_{1}(x_{0})}{1-\beta}P<\frac{1-h_{2}(x_{0})}{\beta}$, i.e., $P_{1}^{\text{active}}<P_{2}^{\text{active}}$.
The same argument applies to the general dependence case by replacing $h_{1}(x_{0})$ with $\sup h_{1}$. Consider the event $U\in[h_{2}(x_{0}),h_{1}(x_{0}))$. In this case, the active $p$-values become
$P_{1}^{\text{active}}=\frac{\sup h_{1}}{1-\beta}P\quad\text{and}\quad P_{2}^{\text{active}}=\frac{1-h_{2}(x_{0})}{\beta}.$
Note that since $h_{1}(x_{0})>h_{2}(x_{0})\geq 1-\beta$, we have $\sup h_{1}>1-\beta$, ensuring that $\frac{1-\beta}{\beta}\cdot\frac{1-h_{2}(x_{0})}{\sup h_{1}}\in(0,1)$. Consequently, with probability $1-\frac{1-\beta}{\beta}\cdot\frac{1-h_{2}(x_{0})}{\sup h_{1}}$, we have
$P>\frac{1-\beta}{\beta}\cdot\frac{1-h_{2}(x_{0})}{\sup h_{1}},$
which is equivalent to $\frac{\sup h_{1}}{1-\beta}P>\frac{1-h_{2}(x_{0})}{\beta}$, i.e., $P_{1}^{\text{active}}>P_{2}^{\text{active}}$. Similarly, with probability $\frac{1-\beta}{\beta}\cdot\frac{1-h_{2}(x_{0})}{\sup h_{1}}$, we have
$P<\frac{1-\beta}{\beta}\cdot\frac{1-h_{2}(x_{0})}{\sup h_{1}},$
which is equivalent to $\frac{\sup h_{1}}{1-\beta}P<\frac{1-h_{2}(x_{0})}{\beta}$, i.e., $P_{1}^{\text{active}}<P_{2}^{\text{active}}$.
Because we have identified events with positive probability where each function yields a strictly better outcome, neither can uniformly dominate the other. Thus, every choice of $h(\cdot)$ with lower bound $1-\beta$ is admissible.
D.5 Proof of Proposition 2
Proof.
We prove the two parts of the proposition—the admissibility of the hyperparameter $\beta$ for the $e$-value setting and $p$-value setting.
Part 1: $e$-value setting
Let $\beta_{1},\beta_{2}\in(0,1)$ be two distinct values, and assume without loss of generality that $\beta_{1}>\beta_{2}$. The proof proceeds by considering two cases based on the range of the control function $h(\cdot)$.
Case 1: $h$ takes an intermediate value.
Assume there exists a point $x_{0}$ such that $0<h(x_{0})<1$. We again consider the setting where $\mathbb{P}(E^{a}=x_{0})=1$. Both branches of the active $e$-value construction are chosen with positive probability.
  • If $U\geq h(x_{0})$, the proxy-based branch is chosen. The resulting $e$-values are $E_{1}^{\text{active}}=\frac{\beta_{1}}{1-h(x_{0})}$ and $E_{2}^{\text{active}}=\frac{\beta_{2}}{1-h(x_{0})}$. Since $\beta_{1}>\beta_{2}$, this immediately yields $E_{1}^{\text{active}}>E_{2}^{\text{active}}$.
  • If $U<h(x_{0})$, the exact-value branch is chosen. The $e$-values are $E_{1}^{\text{active}}=\frac{1-\beta_{1}}{h(x_{0})}E$ and $E_{2}^{\text{active}}=\frac{1-\beta_{2}}{h(x_{0})}E$. Since $\beta_{1}>\beta_{2}$, we have $1-\beta_{1}<1-\beta_{2}$, which for any $E>0$ implies $E_{1}^{\text{active}}<E_{2}^{\text{active}}$.
As both outcomes occur with positive probability, neither choice of $\beta$ dominates the other.
Case 2: $h$ takes only binary values $\{0,1\}$.
Now consider the case where $h$ is non-constant but its range is restricted to $\{0,1\}$. There must exist points $x_{0},x_{1}$ such that $h(x_{0})=0$ and $h(x_{1})=1$. We construct two different distributions for $E^{a}$ to show that neither $\beta_{1}$ nor $\beta_{2}$ can uniformly dominate.
  • Let $\mathbb{P}(E^{a}=x_{0})=1$, where $h(x_{0})=0$. The exact $e$-value branch is never chosen ($U<0$ is impossible). The active $e$-value is always determined by the proxy branch, yielding the deterministic outcomes $E_{1}^{\text{active}}=\beta_{1}$ and $E_{2}^{\text{active}}=\beta_{2}$. As $\beta_{1}>\beta_{2}$, the construction with $\beta_{1}$ is strictly superior in this scenario.
  • Let $\mathbb{P}(E^{a}=x_{1})=1$, where $h(x_{1})=1$. The proxy branch is never chosen ($U\geq 1$ is a zero-probability event). The active $e$-value is always determined by the exact branch, yielding $E_{1}^{\text{active}}=(1-\beta_{1})E$ and $E_{2}^{\text{active}}=(1-\beta_{2})E$. As $\beta_{1}>\beta_{2}$, it follows that $1-\beta_{1}<1-\beta_{2}$, making the construction with $\beta_{2}$ strictly superior for any $E>0$.
Since we have constructed scenarios where each choice of $\beta$ is strictly better, neither can dominate the other. This completes the proof of admissibility for all non-trivial choices of $h$ and $\beta\in(0,1)$.
Part 2: $p$-value setting.
Let $\beta_{1},\beta_{2}\in(0,1)$ be two distinct values, assuming without loss of generality that $\beta_{1}>\beta_{2}$. We proceed by considering two cases based on the range of $h(\cdot)$.
Case 1: $h$ takes an intermediate value.
Assume there exists a point $x_{0}$ such that $0<h(x_{0})<1$. We analyze the setting where $\mathbb{P}(P^{a}=x_{0})=1$.
  • If $U\geq h(x_{0})$, the proxy-based branch is chosen. The resulting $p$-values are $P_{1}^{\text{active}}=\frac{1-h(x_{0})}{\beta_{1}}$ and $P_{2}^{\text{active}}=\frac{1-h(x_{0})}{\beta_{2}}$. Since $\beta_{1}>\beta_{2}$, this yields $P_{1}^{\text{active}}<P_{2}^{\text{active}}$. Here, $\beta_{1}$ is strictly better.
  • If $U<h(x_{0})$, the exact-value branch is chosen. The $p$-values are $P_{1}^{\text{active}}=\frac{C\cdot P}{1-\beta_{1}}$ and $P_{2}^{\text{active}}=\frac{C\cdot P}{1-\beta_{2}}$, where $C$ is a positive constant independent of $\beta$. Since $\beta_{1}>\beta_{2}$, we have $1-\beta_{1}<1-\beta_{2}$, which implies $P_{1}^{\text{active}}>P_{2}^{\text{active}}$ for any $P>0$. Here, $\beta_{2}$ is strictly better.
Since each choice of $\beta$ is strictly better on events with positive probability, neither can dominate.
Case 2: $h$ takes only binary values $\{0,1\}$.
Assume $h$ is non-constant, so there exist points $x_{0},x_{1}$ with $h(x_{0})=0$ and $h(x_{1})=1$.
  • Let $\mathbb{P}(P^{a}=x_{0})=1$. The active $p$-value is always determined by the proxy branch, yielding the deterministic outcomes $P_{1}^{\text{active}}=1/\beta_{1}$ and $P_{2}^{\text{active}}=1/\beta_{2}$. As $\beta_{1}>\beta_{2}$, $P_{1}^{\text{active}}<P_{2}^{\text{active}}$, making the construction with $\beta_{1}$ strictly better.
  • Let $\mathbb{P}(P^{a}=x_{1})=1$. The active $p$-value is always determined by the exact-value branch. This yields $P_{1}^{\text{active}}=C/(1-\beta_{1})$ and $P_{2}^{\text{active}}=C/(1-\beta_{2})$, where $C=1$ in both dependence cases. As $\beta_{1}>\beta_{2}$, we have $1-\beta_{1}<1-\beta_{2}$, which implies $P_{1}^{\text{active}}>P_{2}^{\text{active}}$. The construction with $\beta_{2}$ is strictly better.
Having constructed scenarios where each choice of $\beta$ is strictly superior, we conclude that no choice can uniformly dominate another. This completes the proof of admissibility. ∎
D.6 Proof of Proposition 3
Proof.
The proof proceeds as follows: first, we show that $C_{i}\in\{0,1\}$; second, we verify $\mathbb{E}[C_{i}]=p_{i}$; and finally, we demonstrate that the exact budget constraint $\sum_{i=1}^{N}C_{i}=n_{b}$ holds.
Support of $C_{i}$.
By definition, $S_{i}-S_{i-1}=p_{i}\in[0,1]$. Let $x=S_{i-1}-U$. We can rewrite the indicator as
$C_{i}=\lfloor x+p_{i}\rfloor-\lfloor x\rfloor.$
Since $0\leq p_{i}\leq 1$, it follows that $x\leq x+p_{i}\leq x+1$. The monotonicity of the floor function implies $\lfloor x\rfloor\leq\lfloor x+p_{i}\rfloor\leq\lfloor x\rfloor+1$. Consequently, $0\leq C_{i}\leq 1$. Because $C_{i}$ is defined as the difference between two integers, it must hold that $C_{i}\in\{0,1\}$.
Expectation $\mathbb{E}[C_{i}]=p_{i}$.
Consider the expectation of $\lfloor c-U\rfloor$ for an arbitrary constant $c\in\mathbb{R}$ and $U\sim\mathrm{Uniform}(0,1)$. We decompose $c$ into its integer and fractional parts: $c=\lfloor c\rfloor+\{c\}$, where $\{c\}\in[0,1)$. The random variable $\lfloor c-U\rfloor$ evaluates to $\lfloor c\rfloor$ if $U\leq\{c\}$, and to $\lfloor c\rfloor-1$ if $U>\{c\}$. Its expectation is therefore
$\mathbb{E}[\lfloor c-U\rfloor]$ $=\lfloor c\rfloor\cdot\mathbb{P}(U\leq\{c\})+(\lfloor c\rfloor-1)\cdot\mathbb{P}(U>\{c\})$
$=\lfloor c\rfloor\{c\}+(\lfloor c\rfloor-1)(1-\{c\})$
$=\lfloor c\rfloor-1+\{c\}$
$=c-1.$
So we have
$\mathbb{E}[C_{i}]=\mathbb{E}[\lfloor S_{i}-U\rfloor]-\mathbb{E}[\lfloor S_{i-1}-U\rfloor]=(S_{i}-1)-(S_{i-1}-1)=S_{i}-S_{i-1}=p_{i}.$
Exact sum constraint.
Summing $C_{i}$ over all NN variables yields:
$\sum_{i=1}^{N}C_{i}=\sum_{i=1}^{N}\left(\lfloor S_{i}-U\rfloor-\lfloor S_{i-1}-U\rfloor\right)=\lfloor S_{N}-U\rfloor-\lfloor S_{0}-U\rfloor.$
By construction, $S_{0}=0$ and $S_{N}=\sum_{j=1}^{N}p_{j}=n_{b}\in\mathbb{N}$. Then we have
$\sum_{i=1}^{N}C_{i}=\lfloor n_{b}-U\rfloor-\lfloor-U\rfloor=n_{b}+\lfloor-U\rfloor-\lfloor-U\rfloor=n_{b}.$
D.7 Proof of Proposition B.1
Proof.
We prove the result separately for the $p$-value and $e$-value settings. In both cases, the proof relies on the contradiction arising from the coupling of hypotheses via the budget constraint.
Part 1: $p$-value setting.
Suppose, for the sake of contradiction, that there exists a distinct control vector $\tilde{\mathbf{h}}=(\tilde{h}_{1},\dots,\tilde{h}_{N})\in\mathcal{H}$ whose induced active $p$-value vector $\mathbf{X}_{\tilde{\mathbf{h}},\boldsymbol{\beta}}^{\mathrm{active}}$ dominates $\mathbf{X}_{\mathbf{h},\boldsymbol{\beta}}^{\mathrm{active}}$.
Domination implies that for every component $i$, the statistic induced by $\tilde{h}_{i}$ must be essentially no worse than that induced by $h_{i}$. We first show that this requirement forbids the case where $h_{i}(\mathbf{x})>\tilde{h}_{i}(\mathbf{x})$.
Assume there exists an input $\mathbf{x}$ and index $i$ such that $h_{i}(\mathbf{x})>\tilde{h}_{i}(\mathbf{x})$. We consider two sub-cases:
  1. 1.
    If $\tilde{h}_{i}(\mathbf{x})\geq 1-\beta$, then $h_{i}(\mathbf{x})>\tilde{h}_{i}(\mathbf{x})\geq 1-\beta$. However, following the logic in the proof of Proposition 1, if both functions satisfy the condition $\geq 1-\beta$ and differ, neither dominates the other. Thus, for domination to hold, they must coincide, which contradicts $h_{i}(\mathbf{x})>\tilde{h}_{i}(\mathbf{x})$.
  2. 2.
    If $\tilde{h}_{i}(\mathbf{x})<1-\beta$, then consider the event where the auxiliary statistic is $P^{a}\equiv\mathbf{x}$ and the exact $p$-value $P_{i}\sim\mathrm{Uniform}(0,1)$ is small. In the proxy branch (defined by $U_{i}\geq\tilde{h}_{i}(\mathbf{x})$), the active $p$-value is $\tilde{P}_{i}^{\mathrm{active}}=(1-\tilde{h}_{i}(\mathbf{x}))/\beta$. Since $\tilde{h}_{i}(\mathbf{x})<1-\beta$, we have $\tilde{P}_{i}^{\mathrm{active}}>1$, rendering it non-informative.
    Now consider the interval $U_{i}\in[\tilde{h}_{i}(\mathbf{x}),\,h_{i}(\mathbf{x}))$. On this event, the active statistic for $\mathbf{h}$ computes the exact $p$-value (scaled by $b_{i}$), while $\tilde{\mathbf{h}}$ returns the non-informative proxy >1>1. Whenever the exact $p$-value $P_{i}$ is sufficiently small (specifically $P_{i}<(1-\beta)/b_{i}$), we have $\min(1,X_{h_{i}}^{\mathrm{active}})<\min(1,X_{\tilde{h}_{i}}^{\mathrm{active}})$. This contradicts the assumption that $\mathbf{X}_{\tilde{\mathbf{h}},\boldsymbol{\beta}}^{\mathrm{active}}$ dominates $\mathbf{X}_{\mathbf{h},\boldsymbol{\beta}}^{\mathrm{active}}$.
Thus, we conclude that $h_{i}(\mathbf{x})\leq\tilde{h}_{i}(\mathbf{x})$ must hold for all $i$ and $\mathbf{x}$.
Finally, we invoke the budget constraint. Both vectors must satisfy $\sum_{j=1}^{N}h_{j}(\mathbf{x})=\sum_{j=1}^{N}\tilde{h}_{j}(\mathbf{x})=n_{b}$. If strict domination occurred, there would exist some $j$ and $\mathbf{x}$ such that $h_{j}(\mathbf{x})\neq\tilde{h}_{j}(\mathbf{x})$. Based on the result above, this would imply $h_{j}(\mathbf{x})<\tilde{h}_{j}(\mathbf{x})$. To preserve the sum, there must exist some other index $k$ such that $h_{k}(\mathbf{x})>\tilde{h}_{k}(\mathbf{x})$. However, we have already proven that $h_{k}(\mathbf{x})>\tilde{h}_{k}(\mathbf{x})$ is impossible under domination. Therefore, we must have $\mathbf{h}=\tilde{\mathbf{h}}$ almost everywhere, and no strictly dominating vector exists.
Part 2: $e$-value setting.
Suppose, for contradiction, that there exists another vector $\tilde{\mathbf{h}}\in\mathcal{H}$ such that $\mathbf{X}_{\tilde{\mathbf{h}},\boldsymbol{\beta}}^{\mathrm{active}}$ dominates $\mathbf{X}_{\mathbf{h},\boldsymbol{\beta}}^{\mathrm{active}}$.
If $\tilde{\mathbf{h}}\neq\mathbf{h}$, the budget constraint $\sum h_{i}=\sum\tilde{h}_{i}$ implies that the vectors must differ in at least two components in opposite directions. Specifically, there must exist an input $\mathbf{x}$ and indices $j,k$ such that $\tilde{h}_{j}(\mathbf{x})>h_{j}(\mathbf{x})$ and $\tilde{h}_{k}(\mathbf{x})<h_{k}(\mathbf{x})$.
Consider the component $k$ where $h_{k}(\mathbf{x})>\tilde{h}_{k}(\mathbf{x})$. The proof of Proposition 1 establishes that for any single hypothesis, a control function with a higher value cannot be dominated by one with a lower value (except in trivial cases). This contradicts the assumption that the vector $\mathbf{X}_{\tilde{\mathbf{h}},\boldsymbol{\beta}}^{\mathrm{active}}$ component-wise dominates $\mathbf{X}_{\mathbf{h},\boldsymbol{\beta}}^{\mathrm{active}}$. Thus, $\mathbf{h}$ is admissible. ∎
D.8 Proof of Proposition B.2
Proof.
Fix a non-degenerate control-function vector $\mathbf{h}$ and take any two distinct hyperparameter vectors $\boldsymbol{\beta}_{1},\boldsymbol{\beta}_{2}\in(0,1)^{N}$. Since they differ, there must exists an index $i$ such that $(\boldsymbol{\beta}_{1})_{i}\neq(\boldsymbol{\beta}_{2})_{i}$.
We invoke Proposition 2, which states that for a fixed, non-trivial control function, no scalar $\beta$ dominates another. The non-degenerate assumption on $\mathbf{h}$ ensures that $h_{i}$ is not identically 0 or 1, satisfying the condition for Proposition 2.
Consequently, because $(\boldsymbol{\beta}_{1})_{i}\neq(\boldsymbol{\beta}_{2})_{i}$, there exists an event with positive probability under which $(\mathbf{X}^{\text{active}}_{\mathbf{h},\tilde{\boldsymbol{\beta}}})_{i}$ is superior (larger $e$-value or smaller $p$-value) to $(\mathbf{X}^{\text{active}}_{\mathbf{h},\boldsymbol{\beta}})_{i}$ and an event with positive probability under which $(\mathbf{X}^{\text{active}}_{\mathbf{h},\tilde{\boldsymbol{\beta}}})_{i}$ is inferior (smaller $e$-value or larger $p$-value) to $(\mathbf{X}^{\text{active}}_{\mathbf{h},\boldsymbol{\beta}})_{i}$ However, for the vector $\mathbf{X}^{\text{active}}_{\mathbf{h},\tilde{\boldsymbol{\beta}}}$ to dominate $\mathbf{X}^{\text{active}}_{\mathbf{h},\boldsymbol{\beta}}$, it must be essentially no worse in every component almost surely. The existence of the events described above proves that component $i$ fails this condition. Thus, $\tilde{\boldsymbol{\beta}}$ cannot dominate $\boldsymbol{\beta}$, and we conclude that $\mathbf{X}^{\text{active}}_{\mathbf{h},\boldsymbol{\beta}}$ is admissible.
D.9 Proof of Lemma 1
Proof.
If there exists $x_{0}$ such that $f(x_{0})\cdot x_{0}>1$, then:
$\int_{0}^{x_{0}}f(x)dx\geq\int_{0}^{x_{0}}f(x_{0})dx=f(x_{0})\cdot x_{0}>1$
This implies:
$\int_{0}^{1}f(x)dx\geq\int_{0}^{x_{0}}f(x)dx>1$
which is a contradiction. ∎
Appendix E Counterexample for the Choice of $b(\cdot)$ in Theorem 3
This section provides a formal counterexample to demonstrate why the point-wise optimal choice for $b(q)$ under the independence assumption, namely $b(q)=h(q)/(1-\beta)$, is not valid under a general dependence structure.
Setup.
We aim to violate condition (5) of the main text. Let us choose the hyperparameters $\beta=1/2$ and $s=1$. The candidate function for $b(q)$ is therefore b(q)=h(q)/(11/2)=2h(q)b(q)=h(q)/(1-1/2)=2h(q). The validity condition that must be satisfied is:
$\mathbb{E}\left[h(P^{a})\cdot\mathbb{I}\{b(P^{a})P\leq s\}\right]\leq(1-\beta)s=0.5.$
Construction of an Adversarial Joint Distribution.
To construct a counterexample, we define a specific joint distribution for $(P,P^{a})$ that creates a challenging dependence structure. Let the distribution of $h(P^{a})$ be:
$h(P^{a})=\begin{cases}0.4&\text{with probability }1/2\\ 1.0&\text{with probability }1/2\end{cases}$
We then define the conditional distribution of the exact $p$-value $P$ to be negatively correlated with the value of $h(P^{a})$:
$P\mid(h(P^{a})=0.4)$ $\sim\mathrm{Uniform}(0.5,1)$
$P\mid(h(P^{a})=1.0)$ $\sim\mathrm{Uniform}(0,0.5)$
A straightforward calculation confirms that the marginal distribution of $P$ is $\mathrm{Uniform}(0,1)$, ensuring it is a valid null $p$-value.
Violation of the Validity Condition.
We now compute the left-hand side of the validity condition under this distribution.
$\mathbb{E}\left[h(P^{a})\cdot\mathbb{I}\{b(P^{a})P\leq 1\}\right]$
$=\frac{1}{2}\cdot\mathbb{E}\left[0.4\cdot\mathbb{I}\{(2\cdot 0.4)P\leq 1\}\mid h(P^{a})=0.4\right]$
$\quad+\frac{1}{2}\cdot\mathbb{E}\left[1.0\cdot\mathbb{I}\{(2\cdot 1.0)P\leq 1\}\mid h(P^{a})=1.0\right]$
$=0.2\cdot\mathbb{P}\left(0.8P\leq 1\mid P\sim\mathrm{Uniform}(0.5,1)\right)$
$\quad+0.5\cdot\mathbb{P}\left(2P\leq 1\mid P\sim\mathrm{Uniform}(0,0.5)\right).$
We evaluate the two conditional probabilities.
  • For the first term, when $P\sim\mathrm{Uniform}(0.5,1)$, the value of $0.8P$ is always in the interval $[0.4,0.8]$. Thus, the condition $0.8P\leq 1$ is always true, and the probability is 1.
  • For the second term, when $P\sim\mathrm{Uniform}(0,0.5)$, the value of $2P$ is always in the interval $[0,1]$. Thus, the condition $2P\leq 1$ is also always true, and this probability is 1.
Substituting these probabilities back into the expectation gives:
$\mathbb{E}\left[h(P^{a})\cdot\mathbb{I}\{b(P^{a})P\leq 1\}\right]=0.2\cdot 1+0.5\cdot 1=0.7.$
Conclusion.
The calculated expectation is $0.7$, while the validity condition requires the expectation to be no greater than $0.5$. Since $0.7\not\leq 0.5$, the condition is violated. This demonstrates that the choice $b(q)=h(q)/(1-\beta)$ is not valid in general and underscores the necessity of the more conservative construction for the case of arbitrary dependence.
Appendix F Counterexample for the Decomposition in Remark 3
In our main analysis, we adopted a decomposition strategy, ensuring the validity of the active $p$-value by separately controlling the two components of its tail probability, as shown in equations (4) and (5). A natural question arises: is this decomposition necessary? That is, for any valid active $p$-value construction satisfying the super-uniformity condition, must there exist a universal $\beta\in[0,1]$ that validates the decomposition?
We show that the answer is no. We construct a simple, valid active $p$-value for which no such universal $\beta$ can be found.
A Valid Construction That Defies Decomposition.
Consider the specific construction where $a(p)\equiv 1$ and $b(p)\equiv 1$ for all $p\in[0,1]$. Let $h:[0,1]\to[0,1]$ be any non-constant function (e.g., $h(x)=x$). The active $p$-value is then:
$P^{\mathrm{active}}=\begin{cases}1&\text{if }U\geq h(P^{a})\\ P&\text{if }U<h(P^{a})\end{cases}$
where $U\sim\mathrm{Uniform}(0,1)$ is independent of $(P,P^{a})$. This construction is demonstrably valid. For any $s\in[0,1)$, the proxy branch can never produce a value $\leq s$. Therefore,
$\mathbb{P}(P^{\mathrm{active}}\leq s)=\mathbb{P}(U<h(P^{a})\text{ and }P\leq s)\leq\mathbb{P}(P\leq s)\leq s.$
The super-uniformity condition holds for any valid $(P,P^{a})$ distribution.
Deriving the Contradiction.
Now, assume for the sake of contradiction that there exists a universal $\beta\in[0,1]$ for which the decomposition conditions (4) and (5) hold for this construction.
First, we analyze condition (4). Since $a(P^{a})\equiv 1$, the indicator $\mathbb{I}\{a(P^{a})\leq s\}$ is 0 for $s<1$ and 1 for $s=1$. The condition is trivially satisfied for $s<1$. For $s=1$, it requires:
$\mathbb{E}[(1-h(P^{a}))\cdot\mathbb{I}\{1\leq 1\}]=\mathbb{E}[1-h(P^{a})]\leq\beta\cdot 1.$
This inequality must hold for any distribution of $P^{a}$. If we choose a deterministic $P^{a}\equiv x$ for any $x\in[0,1]$, this implies $1-h(x)\leq\beta$, which rearranges to a lower bound on $h(x)$:
$h(x)\geq 1-\beta\quad\text{for all }x\in[0,1].$ (F.1)
Next, we analyze condition (5). With $b(P^{a})\equiv 1$, it states:
$\mathbb{E}[h(P^{a})\cdot\mathbb{I}\{P\leq s\}]\leq(1-\beta)s.$
To isolate $h(x)$, we can again choose a deterministic $P^{a}\equiv x$ and an independent $P\sim\mathrm{Uniform}(0,1)$. The condition becomes:
$h(x)\cdot\mathbb{P}(P\leq s)=h(x)\cdot s\leq(1-\beta)s.$
This must hold for all $s\in(0,1]$, which implies an upper bound on $h(x)$:
$h(x)\leq 1-\beta\quad\text{for all }x\in[0,1].$ (F.2)
Combining the lower bound from (F.1) and the upper bound from (F.2), we find that for a universal $\beta$ to exist, the function $h(x)$ must satisfy $h(x)=1-\beta$ for all $x\in[0,1]$. This means $h(x)$ must be a constant function.
This contradicts our initial premise that $h(x)$ is a non-constant function. Therefore, our assumption that a universal $\beta$ exists must be false. This counterexample confirms that the decomposition is a sufficient, but not necessary, condition for the validity of an active $p$-value.
Appendix G Comparison with the Framework of Xu et al. (2025b)
In this section, we formalize the relationship between our active testing framework and the closely related method of Xu et al. (2025b). We show that for $e$-values, our construction provides a point-wise dominant statistic, yielding greater power for an identical computational cost. For $p$-values, our construction is strictly more powerful under independence, while under general dependence, the two frameworks are equivalent, revealing the Xu et al. (2025b) construction to be a special case of ours.
G.1 Comparison of $e$-value Constructions
We begin by comparing the active $e$-value constructions. The Xu et al. (2025b) method defines a query probability based on an auxiliary $e$-value $E^{a}$ and a hyperparameter $\beta\in(0,1)$. A Bernoulli random variable $T\sim\mathrm{Bern}((1-\beta(E^{a})^{-1})_{+})$ determines whether to query the exact $e$-value $E$. The final statistic is reported as:
[…]
To establish a direct comparison, we adopt the identical decision rule in our framework by setting the control function to $h(x)=(1-\beta x^{-1})_{+}$. Our active $e$-value is then constructed as:
$E^{\mathrm{active}}:=\begin{cases}\max\{\beta,E^{a}\}&\text{if }T=0\\ (1-\beta)\dfrac{E^{a}}{E^{a}-\beta}E&\text{if }T=1\end{cases}.\quad\text{(Our construction)}$
Derivation of the Proxy-Branch Term.
The $\max\{\beta,E^{a}\}$ term in our construction for the $T=0$ (proxy) branch arises directly from the optimal form of an active $e$-value given in Corollary 1, which is $\beta/(1-h(E^{a}))$. With our specific choice of $h(E^{a})=(1-\beta(E^{a})^{-1})_{+}$, we analyze the denominator in two cases:
  • If $E^{a}\leq\beta$, then $1-\beta(E^{a})^{-1}\leq 0$, so $h(E^{a})=0$. The term becomes $\beta/(1-0)=\beta$.
  • If $E^{a}>\beta$, then $h(E^{a})=1-\beta(E^{a})^{-1}$. The term becomes $\beta/\left(1-(1-\beta(E^{a})^{-1})\right)=\beta/\left(\beta(E^{a})^{-1}\right)=E^{a}$.
Combining these two cases, where the result is $\beta$ if $E^{a}\leq\beta$ and $E^{a}$ if $E^{a}>\beta$, gives precisely $\max\{\beta,E^{a}\}$.
Case 1: $E^{a}<\beta$.
The query probability is $(1-\beta/E^{a})_{+}=0$, so $T=0$ almost surely. The resulting statistics are deterministic:
$\tilde{E}=E^{a}\quad\text{and}\quad E^{\mathrm{active}}=\beta.$
Since $E^{a}<\beta$, our construction yields a strictly larger $e$-value, $E^{\mathrm{active}}>\tilde{E}$.
Case 2: $E^{a}\geq\beta$.
Both outcomes for $T$ occur with positive probability. Conditional on $T$, the statistics are:
$\tilde{E}=\begin{cases}E^{a}&\text{if }T=0\\ (1-\beta)E&\text{if }T=1\end{cases}\quad\text{and}\quad E^{\mathrm{active}}=\begin{cases}E^{a}&\text{if }T=0\\ (1-\beta)\dfrac{E^{a}}{E^{a}-\beta}E&\text{if }T=1\end{cases}.$
When $T=1$, the scaling factor in our construction satisfies $\frac{E^{a}}{E^{a}-\beta}\geq 1$ (with strict inequality for $E^{a}>\beta$). This implies $E^{\mathrm{active}}\geq\tilde{E}$ on the event $\{T=1\}$.
In summary, our construction dominates that of Xu et al. (2025b) point-wise:
  • Almost sure inequality: $E^{\mathrm{active}}\geq\tilde{E}$.
  • Strict improvement: The inequality is strict whenever $E^{a}<\beta$. When $E^{a}>\beta$, it is strict on the event $\{T=1\}$, which occurs with positive probability.
G.2 Comparison of $p$-value Constructions
Next, we compare the active $p$-value constructions. The Xu et al. (2025b) method uses a query probability based on an auxiliary $p$-value $P^{a}$ and defines $T\sim\mathrm{Bern}((1-\beta P^{a})_{+})$. The final statistic is:
[…]
We adopt the same decision rule by setting our control function $h(x)=(1-\beta x)_{+}$ and letting $T:=\mathbb{I}\{U<h(P^{a})\}$ for $U\sim\mathrm{Uniform}(0,1)$.
Independent Setting.
Our active $p$-value under independence is given by:
$P^{\mathrm{active}}=(1-T)\frac{1-h(P^{a})}{\beta}+T\frac{h(P^{a})}{1-\beta}P.$
We compare this to $\tilde{P}$ in each branch of the random trial $T$.
  • Conditional on $T=0$: $\tilde{P}=P^{a}$. Our construction yields $P^{\mathrm{active}}=\frac{1-h(P^{a})}{\beta}=\frac{1-\max\{0,1-\beta P^{a}\}}{\beta}=\min\{\beta^{-1},P^{a}\}$. Since $\beta\in(0,1)$, $\beta^{-1}>1$, and since $P^{a}\in[0,1]$, it follows that $\min\{\beta^{-1},P^{a}\}=P^{a}$. Thus, $P^{\mathrm{active}}=\tilde{P}$.
  • Conditional on $T=1$: $\tilde{P}=\frac{P}{1-\beta}$. Our construction yields $P^{\mathrm{active}}=\frac{h(P^{a})}{1-\beta}P=\frac{\max\{0,1-\beta P^{a}\}}{1-\beta}P$. Since $\max\{0,1-\beta P^{a}\}\leq 1$, we have $P^{\mathrm{active}}\leq\tilde{P}$, with strict inequality whenever $P^{a}>0$ and $P>0$.
Because the statistics are identical in one branch and ours is strictly smaller in the other, our construction is point-wise smaller and thus strictly more powerful under independence.
General Dependence Setting.
Our active $p$-value under general dependence takes the form:
$P^{\mathrm{active}}=(1-T)\frac{1-h(P^{a})}{\beta}+T\frac{1}{1-\beta}P.$
As shown above, the term for the $T=0$ branch simplifies to $P^{a}$. The term for the $T=1$ branch is identical to that of $\tilde{P}$. The entire expression is therefore:
$P^{\mathrm{active}}=(1-T)P^{a}+T(1-\beta)^{-1}P=\tilde{P}$.
The two constructions are identical. This reveals that the Xu et al. (2025b) procedure arises as a special case of our more general framework when the conservative construction for arbitrary dependence is employed.
Appendix H Extension to the Online Setting
While the primary focus of this paper is the batch setting where all auxiliary statistics $\{X_{i}^{a}\}_{i=1}^{N}$ are available simultaneously, our framework can be naturally adapted to an online sequence where hypotheses arrive one by one over time $t=1,2,\dots$. This extension is particularly valuable when the total number of hypotheses NN is unknown or potentially infinite, a common scenario in streaming data applications.
The key challenge in the online setting is the budget management. In the batch setting, we can guarantee exact budget adherence. In contrast, an online procedure must make irrevocable decisions without knowledge of future hypotheses, creating a risk of either premature budget exhaustion or underutilization. Our goal is to design an adaptive allocation strategy that spreads the budget appropriately over time while still prioritizing promising hypotheses.
The theoretical foundation for this extension remains unchanged: the validity of active statistics requires only that the control value $h_{t}$ forms a predictable process. In other words, $h_{t}$ may depend on the historical filtration $\mathcal{F}_{t-1}$ and the current auxiliary statistic $X_{t}^{a}$, but not on future information. Let $\mathcal{B}_{t-1}=n_{b}-S_{t-1}$ denote the remaining budget at time $t$, where $S_{t-1}=\sum_{i=1}^{t-1}C_{i}$ is the cumulative number of expensive tests already performed. We propose the following adaptive allocation rule:
$h_{t}=\min\left(1,\quad\underbrace{\frac{\mathcal{B}_{t-1}\cdot a_{t}}{A_{t}}}_{\text{baseline pacing}}\cdot\underbrace{\left(\frac{u_{t}}{\bar{u}_{t-1}}\right)}_{\text{signal adjustment}}\cdot\underbrace{\exp\left(\eta\cdot\Delta_{t}\right)}_{\text{feedback control}}\right),$ (H.1)
where $u_{t}$ is the base utility of the $t$-th hypothesis, $\bar{u}_{t-1}=\frac{1}{t-1}\sum_{i=1}^{t-1}u_{i}$ is the empirical mean of historical utilities, and $\{a_{t}\}_{t=1}^{\infty}$ is a pre-specified positive sequence with $\sum_{t=1}^{\infty}a_{t}=1$, with $A_{t}=\sum_{j=t}^{\infty}a_{j}$ denoting the remaining mass. Each component of (H.1) serves a distinct purpose:
Baseline pacing. The term $\mathcal{B}_{t-1}\cdot a_{t}/A_{t}$ allocates the remaining budget according to a pre-specified schedule. Intuitively, $a_{t}/A_{t}$ represents the fraction of remaining budget that should be allocated at time $t$ under the nominal schedule. This ensures the budget stretches indefinitely without expiring prematurely, even when NN is unknown.
Signal adjustment. The factor $u_{t}/(\bar{u}_{t-1})$ dynamically adjusts the allocation based on the relative promise of the current hypothesis. When $u_{t}$ exceeds the historical average $\bar{u}_{t-1}$, the allocation probability is boosted, prioritizing hypotheses with stronger auxiliary signals.
Feedback control. The term $\exp(\eta\cdot\Delta_{t})$ acts as a stabilizing mechanism that corrects for deviations from the planned spending trajectory. Let $L_{t}=n_{b}\cdot(1-A_{t+1})$ denote the cumulative budget that should have been consumed by time $t$ under the nominal schedule $\{a_{t}\}$. The deviation $\Delta_{t}=L_{t}-S_{t-1}$ measures whether actual spending $S_{t-1}$ is ahead of or behind schedule. When $\Delta_{t}<0$ (overspending), the exponential term decreases subsequent allocation probabilities; when $\Delta_{t}>0$ (underspending), it increases them. The parameter $\eta>0$ controls the strength of this feedback.
This design naturally satisfies both the budget constraint and statistical validity. If the budget is exhausted ($\mathcal{B}_{t-1}=0$), then $h_{t}=0$ and no further queries are made. Furthermore, because $h_{t}$ uses only past information $\mathcal{F}_{t-1}$ and the current proxy $X_{t}^{a}$, it is a predictable process that guarantees valid active statistics.